### v2

# Perfect Matching in Product Graphs and in their Random Subgraphs

Abstract
For $t\in\mathbb{N}$ and every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph, with $1<|V(H_{i})|\leq C$ for some integer $C\geq 2$. Let $G=\square_{i=1}^{t}H_{i}$ be the Cartesian product of $H_{1},\ldots,H_{t}$. We show that if $t\geq 5C\log_{2}C$ then $G$ contains a (nearly-)perfect matching.
Then, considering the random graph process on $G$, we generalise the result of Bollobás on the binary hypercube $Q^{t}$, showing that with high probability, the hitting times for minimum degree one, connectivity, and the existence of a (nearly-)perfect matching in the $G$-random-process are the same. We develop several tools which may be of independent interest in a more general setting of the typical existence of a perfect matching under percolation.
1 Introduction
1.1 Background and main results
Given two graphs $H_{1}=(V_{1},E_{1})$ and $H_{2}=(V_{2},E_{2})$, the Cartesian product $H=H_{1}\square H_{2}$ is the graph whose vertex set is $V_{1}\times V_{2}$, and for $u_{1},v_{1}\in V_{1}$ and $u_{2},v_{2}\in V_{2}$, we have that $\{(u_{1},u_{2}),(v_{1},v_{2})\}\in E(H)$ either if $u_{1}=v_{1}$ and $\{u_{2},v_{2}\}\in E(H_{2})$ or if $u_{2}=v_{2}$ and $\{u_{1},v_{1}\}\in E(H_{1})$. More generally, given $t$ graphs, $H_{i},\ldots,H_{t}$, their Cartesian product $G=\square_{i=1}^{t}H_{i}$ is the graph with the vertex set
$V\coloneqq\left\{v=(v_{1},\ldots,v_{t})\colon v_{i}\in V(H_{i})\text{ for all }i\in[t]\right\},$
and the edge set
$\left\{uv\colon\begin{array}[]{l}\text{there is some }i\in[t]\text{ such that }u_{j}=v_{j}\\ \text{for all }i\neq j\text{ and }\{u_{i},v_{i}\}\in E\left(H_{i}\right)\end{array}\right\}.$
We call $H_{1},H_{2},\ldots,H_{t}$ the base graphs of $G$. Cartesian product graphs arise naturally in many contexts and have received much attention in combinatorics, probability, and computer science. Many classical graphs, which have been extensively studied, are in fact Cartesian product graphs: the $t$-dimensional torus is the Cartesian product of $t$ copies of the cycle $C_{k}$, the $t$-dimensional grid is the Cartesian product of $t$ copies of the path $P_{k}$, and the binary $t$-dimensional hypercube $Q^{t}$ is the Cartesian product of $t$ copies of a single edge $K_{2}$. We refer the reader to [16] for a systematic coverage of Cartesian product graphs, and related product structures on graphs. Throughout the paper, whenever we write product graphs we consider the Cartesian product as defined above.
In this paper, we study perfect matchings in product graphs. One can convince oneself that if one of the base graphs has a perfect matching, then the product graph $G$ has a perfect matching as well. Indeed, Kotzig [19] showed something even stronger: let $G=\square_{i=1}^{t}H_{i}$ be such that each $H_{i}$ is regular, then if at least one of the base graphs has a $1$-factorisation (recall that a $1$-factorisation is a decomposition of the edges of the graph into disjoint perfect matchings), or at least two base graphs have perfect matchings, then $G$ has a $1$-factorisation; and furthermore these sufficient conditions are not necessary. However, when none of the base graphs have a perfect matching, it is not clear whether their product graph will have a perfect matching. Indeed, not every connected regular graph has a perfect matching (or nearly-perfect matching, that is, a matching covering all but one vertex when the order of the graph is odd). However, note that given two graphs $H_{1}$ and $H_{2}$, the proportion of vertices covered by a largest matching is at least as large as this proportion in any of the base graphs. Thus, taking the product of many graphs, one can hope to improve this proportion. Indeed, our first result gives a simple sufficient condition for the existence of a perfect matching in product graphs in this spirit.
Theorem 1.
Let $C>1$ be a constant, and let $t\geq 5C\log_{2}C$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$ and suppose that $|V(G)|$ is even. Then $G$ has a perfect matching.
Observe that since each $H_{i}$ is a $d_{i}$-regular connected graph, $G=\square_{i=1}^{t}H_{i}$ is a connected $d$-regular graph with $d=\sum_{i=1}^{t}d_{i}$, and in particular $d=\Theta_{C}(\log|V(G)|)$. Furthermore, we note that we have not tried to optimise the constants in the above. Theorem 1 shows that if the dimension of the product graph is sufficiently large, with respect to the maximum size of the base graphs, then the product graph has a perfect matching — regardless of whether the base graphs themselves contain any perfect matchings. Furthermore, from the same proof it follows that if $G$ has an odd number of vertices, then it has a nearly-perfect matching.
As mentioned above, a well-studied product graph is the binary hypercube, $Q^{t}$, in particular in terms of bond percolation on it. Given a graph $G$ and probability $p$, we form the percolated random subgraph $G_{p}\subseteq G$ by including every edge of $G$ independently with probability $p$ (note that $G(n,p)$ is then $(K_{n})_{p}$). The study of $Q^{t}_{p}$ has been initiated by Sapoženko [22] and by Burtin [8], who showed that the sharp threshold for connectivity is $p^{*}=\frac{1}{2}$: when $p<\frac{1}{2}$, with high probability (whp), that is, with probability tending to one as $t$ tends to infinity, $Q^{t}_{p}$ is disconnected, whereas for $p>\frac{1}{2}$, whp $Q^{t}_{p}$ is connected. Erdős and Spencer [14] conjectured that $Q^{t}_{p}$ undergoes a phase transition with respect to its component structure, that is, the typical emergence of a giant component (a connected component containing a linear fraction of the vertices) around $p=\frac{1}{t}$, similar to that of $G(n,p)$ around $p=\frac{1}{n}$. This conjecture was confirmed by Ajtai, Komlós, and Szemerédi [1], with subsequent work by Bollobás, Kohayakawa, and Łuczak [6]. We refer the reader to [20] for a modern short proof of this result.
In recent years, there has been an effort to generalise these results to a wider family of product graphs. Lichev [21] gave sufficient conditions, in terms of the base graphs, for the typical emergence of a giant component in bond percolation on high-dimensional product graphs. Diskin, Erde, Kang, and Krivelevich [12] improved upon this, giving sufficient and tight conditions for the typical emergence of a giant component. Furthermore, they showed that assuming the base graphs are regular, one can give a rather precise description, similar to that in $G(n,p)$, both of the typical component structure [10], and of the asymptotic combinatorial properties of the giant component [11].
Returning to the dense regime, that is, when $p$ is constant, connectivity and the existence of a perfect matching in $Q^{t}_{p}$ have been studied in detail. In particular, Bollobás [4] obtained a hitting time result for the random graph process on $Q^{t}$. Given a graph $\Gamma$, the random graph process on $\Gamma$ is defined as a random sequence of nested graphs $\Gamma(0)\subseteq\ldots\subseteq\Gamma(|E(\Gamma)|)$ together with an ordering $\sigma$ on $E(\Gamma)$, chosen uniformly at random from among all $|E(\Gamma)|!$ such orderings. We set $\Gamma(0)$ to be the empty graph on $V(\Gamma)$. Given $\Gamma(i)$, with $0\leq i<|E(\Gamma)|$, we form $\Gamma(i+1)$ by adding the $(i+1)$-th edge, according to the ordering $\sigma$, to $\Gamma(i)$. The hitting time of a monotone increasing, non-empty graph property $\mathcal{P}$, is a random variable equal to the index $\tau$ for which $\Gamma(\tau)\in\mathcal{P}$, but $\Gamma(\tau-1)\notin\mathcal{P}$. Note that having minimum degree one, connectivity, and the existence of a perfect matching are all monotone increasing properties. Furthermore, observe that for a graph $\Gamma$ to be connected or to contain a perfect matching, the minimum degree of $\Gamma$ has to be at least one. A classical result of Erdős and Rényi [13], and of Bollobás and Thomason [7], is that in the random process on $K_{n}$, whp the hitting time for minimum degree one, connectivity, and the existence of a perfect matching are the same. In 1990, Bollobás [4] showed that the same phenomenon holds in the random graph process on $Q^{t}$.
Theorem 1.1.
Consider the random graph process on $Q^{t}$. Let $\tau_{1}$ be the hitting time of minimum degree one, let $\tau_{2}$ be the hitting time for connectivity, and let $\tau_{3}$ be the hitting time for the existence of a perfect matching. Then, whp, $\tau_{1}=\tau_{2}=\tau_{3}$.
Subsequent work by Joos determined the threshold for connectivity for Cartesian powers of graphs [17, 18], that is, $\square_{i=1}^{t}H_{i}$ where all the base graphs $H_{i}$ are the same.
Our main result generalises the result of Bollobás [4] to a wider family of product graphs.
Theorem 2.
Let $C>1$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph, with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$, and suppose that $|V(G)|$ is even. Consider the random graph process on $G$. Let $\tau_{1}$ be the hitting time of minimum degree one, let $\tau_{2}$ be the hitting time for connectivity, and let $\tau_{3}$ be the hitting time for the existence of a perfect matching. Then, whp, $\tau_{1}=\tau_{2}=\tau_{3}$.
In fact, we prove the following description of $G_{p}$, from which standard results (see, for example, [5, Theorem 7.4] and [4, Lemma 2]), allow one to derive Theorem 2.
Theorem 3.
Let $C>1$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$, and suppose that $n\coloneqq|V(G)|$ is even. Let $d\coloneqq d(G)=\sum_{i=1}^{t}d_{i}$. Let $\epsilon\geq 0$ be a sufficiently small constant, and let $p$ be such that $(1-p)^{d}=n^{-(1-\epsilon)}$. Then, whp, the following holds in $G_{p}$.
  1. (a)
    There exists a unique giant component, spanning all but $o(n)$ of the vertices. All the other components of $G_{p}$, if there are any, are isolated vertices. Furthermore, every two isolated vertices in $G_{p}$ are at distance at least two in $G$.
  2. (b)
    The giant component of $G_{p}$ has a (nearly-)perfect matching.
A few comments are in place. Note that the threshold probability $p^{*}$, at which the expected number of isolated vertices is zero, is $p^{*}=1-\left(\frac{1}{n}\right)^{1/d}$. Noting that $t\leq d\leq(C-1)t$, we have that $d=\Theta_{C-1}(\log n)$. Thus, $\left(\frac{1}{n}\right)^{1/d}$ is bounded away from zero, and bounded away from one, and in turn, $p^{*}$ is bounded away both from zero and from one. Further, if $p$ is such that $(1-p)^{d}=n^{-(1-\epsilon)}$ for small enough constant $\epsilon\geq 0$, then $p$ is bounded away both from zero and from one.
Furthermore, let us remark that Theorem 3a implies that whp adding any edge to $G_{p}$, which is not contained in the giant component, must connect an isolated vertex to the giant component, and indeed this suffices to show that the hitting times of minimum degree one and connectivity are whp the same. Theorem 3b then shows that whp the only obstacle to a perfect matching in $G_{p}$ itself is the existence of isolated vertices. Moreover, the proof of Theorem 3a does not rely on the product structure of the graph, and only uses the assumptions that the graph is $d$-regular, has optimal edge-expansion for sets of size polynomial in $d$, and has some mild edge-expansion for larger sets. Only in the proof of Theorem 3b we use the product structure explicitly.
Let us briefly remark that since the hypercube is bipartite, in order to find a perfect matching it suffices to apply Hall’s theorem. In our setting, however, the graph $G$ is not necessarily bipartite, which means that in order to show the typical existence of a perfect matching, one needs to consider the Tutte-Berge formula. This requires a much more delicate treatment, which is detailed in the proof outline in the subsequent section, and also requires us to develop tools which are valid for more general graphs (see Lemmas 3.2 and 3.7) and could be of independent interest when treating perfect matchings in the setting of percolation.
The paper is structured as follows. In Section 1.2 we give an outline of the proofs. In Section 2 we prove Theorem 3a, and in Section 3 we prove Theorem 1 and Theorem 3b. Finally, in Section 4 we conclude with a short discussion of the results and possible avenues for future research.
1.2 Proofs outline
Notation and definitions.
Throughout the paper we let $C>1$ be an integer, $(H_{i})_{i=1}^{t}$ be a sequence of $d_{i}$-regular and connected graphs with $1<|V(H_{i})|\leq C$, and let $G=\square_{i=1}^{t}H_{i}$ be their product. We let $G=(V,E)$. We call $t$ the dimension of $G$, and given $u=(u_{1},\ldots,u_{t})\in V$ we call $u_{i}$ the $i$-th coordinate of $u$. Furthermore, we denote by $n:=|V|$ the order of $G$, and by $d\coloneqq d(G)$ the degree of $G$, noting once again that $d=\sum_{i=1}^{t}d_{i}=\Theta_{C}(\log n)$.
Given a graph $\Gamma$ and subsets $S_{1},S_{2}\subseteq V(\Gamma)$ with $S_{1}\cap S_{2}=\varnothing$, we denote by $E_{\Gamma}(S_{1})$ the set of edges in $\Gamma$ whose both endpoints are in $S_{1}$, and by $E_{\Gamma}(S_{1},S_{2})$ the edges in $\Gamma$ with one endpoint in $S_{1}$ and the other endpoint in $S_{2}$. We set $e_{\Gamma}(S_{1})\coloneqq|E_{\Gamma}(S_{1})|$ and $e_{\Gamma}(S_{1},S_{2})\coloneqq|E_{\Gamma}(S_{1},S_{2})|$. If the choice of $\Gamma$ is clear, we may omit the subscript. Moreover, given $S\subseteq V(\Gamma)$, we define $S^{C}$ to be $V(\Gamma)\setminus S$.
Finally, throughout the paper, we let $\epsilon\geq 0$ be a sufficiently small constant, and let $p$ be such that $(1-p)^{d}=n^{-(1-\epsilon)}$, recalling that this implies that $p=\Theta_{C}(1)$.
External results.
We make extensive use of the following generalisations of Harper’s inequality to regular high-dimensional product graphs.
Theorem 1.2 (Theorems 1 and 2 in [11]).
For every $S\subseteq V$ with $1\leq|S|\leq n$,
$e(S,S^{C})\geq\frac{|S|\log_{C}\left(\frac{n}{|S|}\right)}{C-1}\text{ and }$
$e(S,S^{C})\geq|S|(d-(C-1)\log_{2}|S|).$
We also utilise the following bound on the number of trees on $k$ vertices in a $d$-regular graph, that are rooted at a fixed vertex.
Lemma 1.3 (Lemma 2 in [3]).
Let $\Gamma$ be a $d$-regular graph, let $k$ be a positive integer and let $v\in V(\Gamma)$. Denote by $t_{k}(v)$ the number of trees on $k$ vertices rooted at $v$ in $\Gamma$. Then $t_{k}(v)\leq(ed)^{k-1}$.
Proof outline of Theorem 3a.
The proof follows from a double-exposure argument, similar in spirit to the classical argument of Ajtai, Komlós, and Szemerédi [1]. Let $p_{2}\coloneqq\frac{1}{d^{2}}$ and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note that $G_{p_{1}}\cup G_{p_{2}}\sim G_{p}$, and $p_{1}\approx p-\frac{1}{d^{2}}$. We first show, using a first-moment calculation, that whp every two vertices $u,v$ which are isolated in $G_{p_{1}}$ are at distance at least two in $G$ (Lemma 2.1). Then, utilising Theorem 1.2 and Lemma 1.3, we show that whp there are no components in $G_{p_{1}}$ whose order is in $\left[2,d^{20}\right]$ (Lemma 2.2). We then turn to show that typically components whose order is at least $d^{20}$ merge after sprinkling with $p_{2}$ — this is fairly standard, and follows the same approach as in [1]. Noting that any vertex which was an isolated vertex in $G_{p_{1}}$, is whp either an isolated vertex after sprinkling with $p_{2}$ or merged into a component which was of size at least $d^{20}$ in $G_{p_{1}}$, we obtain that whp $G_{p}$ has a unique large connected component, and all the other vertices are isolated vertices. Finally, from Markov’s inequality it follows that the total volume of isolated vertices in $G_{p}$ is at most $n^{2\epsilon}=o(n)$.
Proof outline of Theorem 3b.
Recall that we seek to show that the giant component of $G_{p}$ has a (nearly-)perfect matching. We know from Theorem 3a that whp every vertex of $G$ which is not in the giant component is an isolated vertex in $G_{p}$. Thus, it suffices to show that whp $G_{p}$ has a perfect matching if there are typically no isolated vertices in $G_{p}$. To that end, we will utilise the Tutte-Berge formula [2], which implies that the number of unmatched vertices in a maximum matching in a graph $\Gamma$ is equal to
$\max_{U\subseteq V(\Gamma)}\left(\mathrm{odd}\left(\Gamma[V(\Gamma)\setminus U]\right)-|U|\right),$
where $\mathrm{odd}\left(\Gamma[V(\Gamma)\setminus U]\right)$ is the number of connected components with an odd number of vertices in $\Gamma[V(\Gamma)\setminus U]$. In particular, a graph $\Gamma$ has a perfect matching if and only if for every subset $U\subseteq V(\Gamma)$, the subgraph $\Gamma[V(\Gamma)\setminus U]$ has at most $|U|$ connected components with an odd number of vertices. Note here that if $\Gamma$ has isolated vertices, then choosing $U=\varnothing$ witnesses that there is no perfect matching in $\Gamma$.
We thus define Tutte-like obstructions — we say that $U\subseteq V$ is an obstruction (for $G_{p}$) if $|U|\geq 1$, and the number of components of size different than two in $G_{p}[V\setminus U]$ is at least $|U|+1$. We say that the size of the obstruction is $|U|$. Note that if $G_{p}$ has no such obstructions, then the only possible obstructions to a perfect matching are isolated vertices, and whp the giant component has a perfect matching. Furthermore, observe that this definition does not capture all the properties coming from the Tutte-Berge formula – indeed, we treat components with an even number of vertices, except two, the same as components with an odd number of vertices – but this crude, yet much-simplified, outlook will suffice for our needs.
We denote the components in $G_{p}[V\setminus U]$ of size different from two by $K_{1},\ldots,K_{\ell}$, where $\ell\geq|U|+1$ if $U$ is an obstruction, and let $K=\bigcup_{i=1}^{\ell}K_{i}$ and $k_{i}\coloneqq|K_{i}|$ for every $i\in[\ell]$. We denote the components of size two in $G_{p}[V\setminus U]$ by $W_{1},\ldots,W_{w/2}$, such that their total volume is $w$.
We say that $U$ is a minimal obstruction, if it is an obstruction with the smallest size of $U$. Note that if a graph has an obstruction, then it must also have a minimal obstruction. We will use the convention that $u\coloneqq|U|$ and $k\coloneqq|K|$. Note that $u+k+w=n$.
We will require a finer description of the components in $K$. Let us partition the vertices in $K$ into three sets, according to the size of the component in $G_{p}[V\setminus U]$ they belong to. Let $\ell_{1}$ be the number of components of size one in $G_{p}[V\setminus U]$, let $\ell_{2}$ be the number of components whose size is in $[3,d^{2}]$ in $G_{p}[V\setminus U]$, and let $\ell_{3}=\ell-\ell_{1}-\ell_{2}$ be the number of components in $G_{p}[V\setminus U]$ that contain more than $d^{2}$ vertices. We further denote the set of vertices in components of $G_{p}[V\setminus U]$ whose size is one by $V_{1}$, those whose size is in $[3,d^{2}]$ by $S$, and those whose size is larger than $d^{2}$ by $B$. We write $s:=|S|$ and $b:=|B|$, and note that $\ell_{1}=|V_{1}|$.
Figure 1: Illustration of an obstruction, with the sets $U$, $V_{1}$, $W$, $S$ and $B$. Note that the only edges in $G_{p}$, which are not induced by $U$ or by components $K_{i}$ in $K$ or $W_{i}$ in $W$, are in $E(U,V_{1}\cup W\cup S\cup B)$.
Note that an obstruction is uniquely determined by the choice of $U$, and that $U$ is not necessarily connected, and thus the number of choices could be $\binom{n}{u}$. As is often the case, a key ingredient will be to efficiently enumerate the number of possible obstructions, and $\binom{n}{u}$ will often be an inefficient bound. We thus prove two claims, which could be of independent interest in arguing about perfect matchings under percolation for general graphs. First, in Lemma 3.2, we show that a minimal obstruction of size $u$ is determined by the choice of $W\cup S\cup B$. We stress that this result holds for any graph $G$, without requirements on its degree or structure. As these sets contain connected components, by Lemma 1.3 there are at most (roughly) $n^{w/2+\ell_{2}+\ell_{3}}(ed)^{s+w+b}$ choices for a minimal obstruction of size $u$ (see Lemma 3.3 for a precise statement). Since $d=\Theta_{C}(\log n)$, for ‘large’ $u$ this will be a much more efficient bound than $\binom{n}{u}$. As it turns out, for certain types of obstructions, neither of these bounds will be efficient enough, and we will require a different approach. Utilising the isoperimetric properties of $G$ (Theorem 1.2), we show in Lemma 3.7 that whp if $U$ is an obstruction of order polynomial in $n$ (that is, $n^{a}$ for some constant $10<a<$ 1), then there is a set $M$ such that $U\subseteq M$, $|M|\leq 10u$, and there are relatively few connected components in $M$ — that is, $U$ must reside in a relatively well-connected set of order proportional to $|U|$. This allows us to more efficiently bound the number of choices of such obstructions, using Lemma 1.3.
With these definitions and tools at hand, we can now describe how to proceed in showing that typically obstructions do not exist. Using 1.2 one can deduce a lower bound of the number of edges in $E(K_{i},K_{i}^{C})$. Note that for $U$ to be an obstruction, $E_{G_{p}}(K_{i},K_{i}^{C})\subseteq E(K_{i},U)$ for all $i\in[\ell]$. Since $G$ is $d$-regular, there are at most $du$ edges touching $U$. So in order to have an obstruction in $G_{p}$, many of the edges in $\bigcup_{i\in\ell}E(K_{i},K_{i}^{C})$ cannot be present in $G_{p}$. Utilising the enumeration given by Lemma 3.7 and this probabilistic bound, we show in Lemma 3.8 that whp there are no obstructions with $1\leq u\leq\frac{n}{d^{C^{3}/p}}$. Then, using the enumeration given by Lemma 3.2 and the aforementioned probabilistic bound, we show that whp there are no obstructions with $b$ not ‘too large’. For larger values of $b$, the aforementioned probabilistic bound no longer suffices. Thus, the final ingredient, similar in spirit to the approach of Bollobás in [4], will be Lemma 3.5, which allows to efficiently bound the number of choices for sets $B$ that have ‘bad expansion’, that is, the number of edges in $E(B,B^{C})$ is small. We note that this is the only place where we use the structure of $G$ as a product graph. Indeed, sets $B$ with bad expansion leave a large ‘fingerprint’ on some coordinates of the product, in particular, if $B$ intersects non-trivially with a projection $G(I)$ of $G$ onto some coordinates $I\subseteq[t]$, then $G(I)$ spans many edges of $E(B,B^{C})$.
Finally, let us briefly comment on the difference between the proof here and the proof for the hypercube, given in [4]. Since the hypercube is a bipartite graph, it suffices there to apply Hall’s theorem, and in particular, one does not need to consider the entirety of the giant component, but only sets of size at most $\frac{n}{2}$ which expand badly. Hence, Harper’s isoperimetric inequality for the hypercube, together with a fingerprint argument given by Bollobás [4], allows one to show that whp there are no obstructions (bar isolated vertices) to a perfect matching. Here, since $G$ is not necessarily bipartite, one needs to consider Tutte-type obstructions, and in particular sets encompassing the entirety of the giant component. This, in turn, requires a much more delicate treatment, and in particular, to consider the structure of obstructions, and the more efficient enumerations given by Lemmas 3.2 and 3.7. This part of the analysis is the key novelty of this work.
2 Connectivity
We begin by showing that isolated vertices in $G_{p}$ are typically at distance at least two in $G$.
Lemma 2.1.
Whp, every two isolated vertices in $G_{p}$ are at distance at least two.
Proof.
Fix an edge $\{u,v\}\in E$. Since $G$ is $d$-regular, the total number of edges meeting $u$ and $v$ is $12d-$ 1. Thus, the probability that $u$ and $v$ are isolated vertices in $G_{p}$ is at most
$(1-p)^{2d-1}\leq(1-p)^{-1}n^{-2(1-\epsilon)}\leq n^{-2(1-2\epsilon)},$
where we used that $(1-p)^{d}=n^{-(1-\epsilon)}$. There are $\frac{dn}{2}$ edges to consider. Thus, by the union bound, the probability that two isolated vertices in $G_{p}$ are at distance one in $G$ is at most $\frac{dn}{2}n^{-2(1-2\epsilon)}=o(1)$. ∎
Let us further show the following ‘gap’ statement, which is an almost-immediate corollary of Theorem 1.2 and Lemma 1.3. We note that $d^{20}$ in the following lemma is chosen rather arbitrarily, and a much tighter bound can be obtained.
Lemma 2.2.
Whp, there are no connected components $K$ in $G_{p}$ with $|V(K)|\eqqcolon k\in[2,d^{20}]$.
Proof.
Fix $k\in[2,d^{20}]$. Let us estimate the probability that there is a connected component $K$ of order $k$ in $G_{p}$. By Theorem 1.2, we have that $e(K,K^{C})\geq k(d-(C-1)\log_{2}k)\geq\frac{9kd}{10}$, where we used the fact that $k\leq d^{20}$. Let $\mathcal{T}_{k}$ be the set of trees of order $k$, and let $\mathcal{A}_{k}$ be the event that there exists a connected component of order $k$ in $G_{p}$. Thus, by the union bound and by Lemma 1.3,
$\mathbb{P}\left[\mathcal{A}_{k}\right]$ $\leq\sum_{T\in\mathcal{T}_{k}}(1-p)^{e(T,T^{C})}\leq n(ed)^{k-1}(1-p)^{\frac{9kd}{10}}\leq n^{1-(1-\epsilon)\frac{9k}{10}}(ed)^{k-1}\leq n^{1-(1-2\epsilon)\frac{9k}{10}}.$
By the union bound over the less than $d^{20}$ possible values of $k$, we have that the probability of an event violating the statement of the lemma is at most
$\sum_{k=2}^{d^{20}}n^{1-(1-2\epsilon)\frac{9k}{10}}\leq 2n^{1-(1-2\epsilon)\frac{18}{10}}\leq n^{-7/10}=o(1).$
We are now ready to prove the key result of this section, that is, Theorem 3a, whose proof will utilise the classical double-exposure argument of [1].
Proposition 2.3.
Whp there is a unique giant component in $G_{p}$ whose order is $n-o(n)$, and all the other components in $G_{p}$ are isolated vertices. Furthermore, every two isolated vertices are at distance at least two from each other.
Proof.
Let $p_{2}=\frac{1}{d^{2}}$, and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note there exists $\epsilon^{\prime}>0$ such that $(1-p_{1})^{d}=n^{-(1-\epsilon^{\prime})}$, and thus we may apply Lemmas 2.1 and 2.2 on $G_{p_{1}}$. Furthermore, observe that $G_{p}$ has the same distribution as $G_{p_{1}}\cup G_{p_{2}}$.
We begin by percolating with probability $p_{1}$. Note that by Lemma 2.2, whp in $G_{p_{1}}$ there are only isolated vertices and components of order at least $d^{20}$. We begin by showing that after sprinkling with $p_{2}$, typically all the components of order at least $d^{20}$ merge. Let $W$ be the set of vertices in components of order at least $d^{20}$ in $G_{p_{1}}$. Let $A\sqcup B=W$ be a partition of $W$ which respects the components of $G_{p_{1}}$, that is, every component is contained either in $A$ or in $B$. We may assume that $|A|\leq|B|$ and let $a\coloneqq|A|$.
Let $A^{\prime}\coloneqq\left(A\cup N_{G}(A)\right)\setminus B$, where $N_{G}(X)$ is the external neighbourhood of $X$ in $G$, and let $B^{\prime}\coloneqq\left(B\cup N_{G}(B)\right)\setminus A^{\prime}$. By Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are at distance at least two from each other in $G$. We continue assuming this holds deterministically. Thus, every isolated vertex in $G_{p_{1}}$ is in the neighbourhood of $A$ or $B$, and we have that $V=A^{\prime}\sqcup B^{\prime}$, that is, $B^{\prime}=V\setminus A^{\prime}$. By Theorem 1.2, we have that $e(A^{\prime},B^{\prime})\geq\frac{a\log_{C}(n/a)}{C-1}\geq\frac{a\log_{C}2}{C-1}$, since $a\leq\frac{n}{2}$. We can extend these edges to $\frac{a\log_{C}2}{C-1}$ paths of length at most $3$ between $A$ and $B$. Every two such paths share at most two edges. Thus, we can trim the set of $\frac{a\log_{C}2}{C-1}$ paths to a set of $\frac{a\log_{C}2}{Cd^{2}}$ edge-disjoint paths between $A$ and $B$.
We now sprinkle with probability $p_{2}$. The probability none of these paths are in $G_{p_{2}}$ is thus at most $(1-p_{2}^{3})^{\frac{a\log_{C}2}{Cd^{2}}}\leq\exp\left\{-\frac{a}{d^{9}}\right\}$. For a union bound note that there are at most $\binom{n/d^{20}}{a/d^{20}}$ possibilities to choose the partition of $W$. Thus, the probability that there is a partition of $W$ such that there is no path between two components of order at least $d^{20}$ is at most
$\sum_{a=d^{20}}^{\frac{n}{2d^{20}}}\binom{n/d^{20}}{a/d^{20}}\exp\left\{-\frac{a}{d^{9}}\right\}$ $\leq\sum_{a=d^{20}}^{n}(en)^{a/d^{20}}\exp\left\{-\frac{a}{d^{9}}\right\}=o(1),$
where we used $\binom{m}{k}\leq(em)^{k}$ and the fact that $\ln n\leq\ln(C^{d})\leq d^{2}$, as well as that $a\geq d^{20}$. Thus, the probability that there is a component of order at least $d^{20}$ which does not merge in $G_{p_{1}}\cup G_{p_{2}}$ tends to zero. That is, all the components of order at least $d^{20}$ in $G_{p_{1}}$ merge after sprinkling with probability $p_{2}$.
Note that by Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are not connected by an edge of $G$. Hence, adding any edge touching an isolated vertex connects it to a component whose order is at least $d^{20}$ in $G_{p_{1}}$, and these components all merge whp. Hence, whp, there exists a unique connected component in $G_{p_{1}}$ whose order is at least $d^{20}$, and all the other components are isolated vertices, whose distance in $G$ is at least two. Finally, let $Y$ be the random variable counting the number of isolated vertices in $G_{p}$. Then, $\mathbb{E}[Y]=n(1-p)^{d}\leq n^{\epsilon}$. Thus, by Markov’s inequality, whp $Y\leq n^{2\epsilon}=o(n)$, completing the proof. ∎
3 Perfect matching
We begin with the proof of Theorem 1. Throughout the section, we assume divisibility by two whenever is necessary, noting that minor modifications will allow us to argue for the existence of nearly-perfect matching otherwise. Let us first show that, by Theorem 1.2, for ‘high enough’ dimension of $G$ every edge-cut of $G$ has at least $d$ edges.
Lemma 3.1.
Suppose that $t\geq 5C\log_{2}C$. Then, we have that $G$ is $d$-edge-connected.
Proof.
It suffices to show that $e(S,S^{C})\geq d$ for all $S\subseteq V$ with $11\leq|S|\leq n-$ 1, and in fact by symmetry, for $S$ of size at most $\frac{n}{2}$. By 1.2,
$e(S,S^{C})\geq|S|(d-(C-1)\log_{2}|S|).$ (1)
Taking a derivative in $|S|$ we obtain, that for all $|S|$ such that $d\geq(C-1)\log_{2}|S|+(C-1)/\ln 2$, the right-hand side of (1) is increasing with $|S|$. Consider $|S|\leq d$. Then,
$d-(C-1)\log_{2}(|S|)\geq d-(C-1)\log_{2}d\geq\frac{1.1C\log_{2}C}{\ln 2}>\frac{C}{\ln 2}\,,$
where we used that $d\geq t\geq 5C\log_{2}C$ and $C\geq 2$. Hence, $|S|(d-(C-1)\log_{2}|S|)$ is increasing with $|S|$ for $|S|\leq d$, and for all $S$ with $|S|\leq d$ we have that $e(S,S^{C})\geq d$, as required. For $d\leq|S|\leq 2^{\frac{d-1}{C-1}}$ we have $d-(C-1)\log_{2}|S|\geq 1$, and thus $e(S,S^{C})$ is at least $d$, as required. Thus, suppose that $2^{\frac{d-1}{C-1}}\leq|S|\leq\frac{n}{2}$. Note that $|S|\leq n/2$, which implies that $\log_{C}(n/|S|)\geq\log_{C}2$. Hence, by 1.2,
$e(S,S^{C})$ $\geq\frac{|S|\log_{C}(|S|/2)}{C-1}\geq\frac{|S|\log_{C}2}{C-1}\geq\frac{2^{\frac{d-1}{C-1}}\log_{C}2}{C-1}\geq\frac{2^{\frac{d-1}{C-1}}}{2C\ln C}\,.$ (2)
We claim that for our choice of $t$, (2) is at least $d$. Note that the $\frac{2^{\frac{d-1}{C-1}}}{2C\ln C}-d$ is increasing as $d$ is increasing, and indeed
$\frac{2^{\frac{5C\log_{2}C-1}{C-1}}}{2C\ln C}-5C\log_{2}C>\frac{C^{3}}{2\ln C}-5C\log_{2}C\geq 0,$
for $C\geq 2$. ∎
The proof of Theorem 1 then follows immediately from Lemma 3.1, since every $d$-regular, $(d-1)$-edge-connected graph has a perfect matching (see [2, Theorem 7, Chapter 18]).
3.1 Threshold for a perfect matching
Recall the definitions of an obstruction and the sets $U,V_{1},W,S$ and $B$ given in Section 1.2, and that $p$ satisfies $(1-p)^{d}=n^{-(1-\epsilon)}$ for a sufficiently small constant $\epsilon\geq 0$. We begin by collecting several lemmas which we will utilise to show that whp there are no obstructions with $u\geq 1$.
3.2 Typical properties of obstructions
We begin with the following Lemma, which will allow us to bound the number of possible obstructions more efficiently.
Lemma 3.2.
There are at most two minimal obstructions, $U$ and $U^{\prime}$, with $|U|=|U^{\prime}|=u\geq 2$ and the same choice of $W\cup S\cup B$.
Proof.
Fix a minimal obstruction $U$ with $u=|U|\geq 2$. We have that $V\setminus(U\cup W\cup S\cup B)=V_{1}$.
Suppose first that there are two sets, $\varnothing\neq A_{1}\subsetneq U$ and $\varnothing\neq A_{2}\subseteq V_{1}$, such that $X\coloneqq(U\setminus A_{1})\cup A_{2}$ is a minimal obstruction as well. By minimality, we have that $|A_{1}|=|A_{2}|$. Since $X$ is also an obstruction and has the same set of components of size strictly larger than one, every $v\in V_{1}\setminus A_{2}$ has that $N(v)\subseteq U\setminus A_{1}$. But then, we have that $U\setminus A_{1}$ is also an obstruction, where $1\leq|U\setminus A_{1}|<u$ — contradicting the minimality of $U$. Indeed, there are $\ell_{1}-|A_{2}|$ components of size $1$ in $G_{p}[V\setminus(U\setminus A_{1})]$, and $\ell_{2}+\ell_{3}$ components of size at least three in $G_{p}[S\cup B]$, thus at least $\ell_{1}+\ell_{2}+\ell_{3}-|A_{2}|=\ell-|A_{2}|\geq u+1-|A_{2}|=|X|+1$ components in total.
Now, suppose towards contradiction that there are three minimal obstructions $U,U^{\prime},U^{\prime\prime}$ with the same choice of $W\cup S\cup B$. Then, by the above, we have $\varnothing\neq U^{\prime},U^{\prime\prime}\subseteq V_{1}$ with $U^{\prime}\neq U^{\prime\prime}$, such that $u=|U^{\prime}|=|U^{\prime\prime}|$. But then, note that there must be some $v\in U^{\prime\prime}\subseteq V_{1}$, with $v\notin U^{\prime}$. Since $N(v)\subseteq U$, $N(v)\subseteq U^{\prime}$, and $U^{\prime}\cap U=\varnothing$, we have that $v$ is an isolated vertex. But then $U^{\prime\prime}\setminus\{v\}$ is also an obstruction of size $u-1\geq 1$ — contradiction. ∎
We will often seek to enumerate the number of minimal obstructions, having fixed $w,\ell_{2},s,\ell_{3},$ and $b$.
Lemma 3.3.
Given $u,w,\ell_{2},s,\ell_{3},$ and $b$, the number of minimal obstructions is at most
$2n^{w/2+\ell_{2}+2\ell_{3}}(ed)^{w/2+2s+b}.$
Proof.
Recall that $W$ is a set of $w/2$ edges. We thus have at most $\binom{nd/2}{w/2}\leq(nd)^{w/2}$ ways to choose $W$. As for $S$, it has $\ell_{2}$ components. Let us denote the sizes of these components by $s_{1},\ldots,s_{\ell_{2}}$, where we have that $3\leq s_{i}\leq d^{2}$ for every $i\in[\ell_{2}]$. Thus there are at most $d^{2\ell_{2}}$ ways to choose $s_{1},\ldots,s_{\ell_{2}}$. We then have at most $\binom{n}{\ell_{2}}\leq n^{\ell_{2}}$ ways to choose roots for some spanning trees of these components. Thus by Lemma 1.3, the number of ways to choose $S$ is at most $d^{2\ell_{2}}n^{\ell_{2}}\prod_{i=1}^{\ell_{2}}(ed)^{s_{i}-1}\leq n^{\ell_{2}}(ed)^{2s}$. Finally, $B$ has $\ell_{3}$ components, and here we use the crude bound of at most $n^{\ell_{3}}$ ways to choose their sizes. Then, similarly to the above, by Lemma 1.3 there are at most $n^{2\ell_{3}}(ed)^{b}$ ways to choose $B$.
By Lemma 3.2, given $W\cup S\cup B$ and $u$ there are at most $2$ minimal obstructions. Thus, given $u,w,\ell_{2},s,\ell_{3},$ and $b$, the number of minimal obstructions is at most $2n^{w/2+\ell_{2}+2\ell_{3}}(ed)^{w/2+2s+b}$. ∎
Noting that there are at most $du$ edges touching $U$ in $G$, let us bound from below the number of edges leaving the components $K_{i}$ for $i\in[\ell]$. To that end, for $m\in[n/2]$, let
$f(m)=\max\left\{m(d-(C-1)\log_{2}m),\frac{m\log_{C}(n/m)}{C-1}\right\},$ (3)
where for $m\in[n/2+1,n]$ we set $f(m)=f(n-m)$.
Lemma 3.4.
Suppose $U\subseteq V$ is an obstruction. Then
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})\geq d(\ell-1)+f(k-\ell+1),$
where $f$ is defined according to (3).
Proof.
By Theorem 1.2, we have that $\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})\geq\sum_{i=1}^{\ell}f(k_{i})$. Further,
$\sum_{i=1}^{\ell}f(k_{i})=\sum_{i=1}^{\ell}\max\left\{k_{i}(d-(C-1)\log_{2}k_{i}),\frac{k_{i}\log_{C}(n/k_{i})}{C-1}\right\}.$
We claim that the minimum of the above sum is obtained with $k_{i}=1$ for all $i\in[\ell-1]$, and $k_{\ell}=k-(\ell-1)$. To that end, consider the discrete functions $g_{1}(x)=x(d-(C-1)\log_{2}x)$ and $g_{2}(x)=\frac{x\log_{C}(n/x)}{C-1}$, supported on $\{1,\ldots,n/2\}$, noting that these functions are concave. Let $g=\max\{g_{1},g_{2}\}$, supported on $\{1,\ldots,n/2\}$. It suffices to show that $g$ is concave. Indeed, considering $\{1,\ldots,n/2\}^{\ell}$, we then have a function $k_{1},\ldots,k_{\ell}\to\sum_{i=1}^{\ell}g(k_{i})$, which is the sum of concave functions, and thus concave in itself, and is defined on a convex set and thus attains its minimum at the extremal points, that is, when $k_{i}=1$ for all $i\in[\ell-1]$ and $k_{\ell}=k-(\ell-1)$. To show that $g$ is concave, it suffices to show that its second derivative is decreasing, that is, that $q(m)=g(m)-g(m-1)$ is decreasing. Observe that there is some minimal integer $m_{0}$ such that for all $m\geq m_{0}$, $g_{2}(m)\geq g_{1}(m)$, and for all $m^{\prime}<m_{0}$, $g_{2}(m^{\prime})\leq g_{1}(m^{\prime})$. Thus, $q$ is decreasing on $\{2,\ldots,m_{0}-1\}$ and on $\{m_{0}+1,\ldots,n/2\}$. All that is then left is to verify that $q(m_{0}+1)\leq q(m_{0})\leq q(m_{0}-1)$, which follows from tedious yet standard calculations, using $t\leq d\leq(C-1)t$ and $2^{t}\leq n\leq C^{t}$. ∎
We will further make use of the following estimate on the number of sets whose size is not too small, yet having a small edge-boundary.
Lemma 3.5.
Let $a\geq\frac{n}{d^{\ln^{2}d}}$. Then, the number of sets $A\subseteq V$ of size $a$ with $e(A,A^{C})<a\ln^{2}d$ is at most $\exp\left\{\frac{2a}{\ln d}\right\}$.
Proof.
Let $\mathcal{F}$ be the family of $A\subseteq V$ satisfying the conditions of the lemma.
For $i\in[t]$ and any $A\subseteq V$, let $E_{i}(A,A^{C})\subseteq E(A,A^{C})$ be the set of edges in $E(A,A^{C})$ corresponding to a change in the $i$-th coordinate, and let $e_{i,A}\coloneqq|E_{i}(A,A^{C})|$. Moreover, given $I\subseteq[t]$, let $e_{I,A}=\sum_{i\in I}e_{i,A}$. We say that $A$ is bad with respect to a set of coordinates $I$, if $e_{I,A}<a\ln^{2}d\cdot\frac{|I|}{t}$. Let $\mathcal{A}_{I}$ be the family of sets $A$ which are bad with respect to some $I\subseteq[t]$. Note that for every fixed $m\in\mathbb{N}$, if $A\in\mathcal{F}$, then there is some $I$ with $|I|=m$ such that $e_{I,A}<a\ln^{2}d\cdot\frac{|I|}{t}$. Thus,
$|\mathcal{F}|\leq\sum_{\begin{subarray}{c}I\subseteq[t]\\ |I|=m\end{subarray}}|\mathcal{A}_{I}|\leq\binom{t}{m}\max_{\begin{subarray}{c}I\subseteq[t]\\ |I|=m\end{subarray}}|\mathcal{A}_{I}|.$
We now set $m=\log_{C}\left(\ln^{5}d\right)$, and turn to estimate $|\mathcal{A}_{I}|$ for any $I\subseteq[t]$ with $|I|=m$.
For such an $I$ and $v\in V$, let $G(I,v)\coloneqq\square_{i\in[t]\setminus I}\{v_{i}\}\square_{i\in I}H_{i}\subseteq G$. Observe that $2^{|I|}\leq|V(G(I,v))|\leq C^{|I|}$, and that for every $v\neq u\in V$, $V(G(I,v))$ and $V(G(I,u))$ are either disjoint or identical. Thus, fixing $I$ with $|I|=m$, there are at most $\frac{n}{2^{m}}$ different subgraphs $G(I,v)$, and their union is $V$. We say that $A$ intersects non-trivially with $G(I,v)$ if $V(G(I,v))\cap A\neq\varnothing$ and $V(G(I,v))\setminus A\neq\varnothing$. By Theorem 1.2 and Lemma 3.1, we have that if $A$ intersects non-trivially with $G(I,v)$, then $G(I,v)$ spans at least $|I|=m$ edges of $E(A,A^{C})$. Thus, if $A\in\mathcal{A}_{I}$, we have that $A$ intersects non-trivially at most $\frac{a\ln^{2}d}{t}$ such subgraphs. Indeed, otherwise, there would be no $I$ with $|I|=m$ and $e_{I,A}<a\ln^{2}d\cdot\frac{m}{t}$. Therefore, a set $A\in\mathcal{A}_{I}$ contains at least
$\frac{a-C^{m}\frac{a\ln^{2}d}{t}}{C^{m}}$
such subgraphs, and at most $C^{m}\frac{a\ln^{2}d}{t}$ other vertices. Recalling $a\geq\frac{n}{d^{\ln^{2}d}}$ we thus obtain that
$|\mathcal{A}_{I}|$ $\leq\binom{\frac{n}{2^{m}}}{\frac{a}{C^{m}}}\binom{n}{C^{m}\frac{a\ln^{2}d}{t}}\leq\left(\frac{enC^{m}}{a}\right)^{\frac{a}{C^{m}}}\left(\frac{ent}{C^{m}a\ln^{2}d}\right)^{C^{m}\frac{a\ln^{2}d}{t}}$
$\leq\left(ed^{\ln^{2}d}C^{m}\right)^{\frac{a}{C^{m}}}\left(\frac{etd^{\ln^{2}d}}{C^{m}\ln^{2}d}\right)^{C^{m}\frac{a\ln^{2}d}{t}}\leq\left(ed^{\ln^{2}d}\ln^{5}d\right)^{\frac{a}{\ln^{5}d}}\left(\frac{ed^{\ln^{2}d+1}}{\ln^{7}d}\right)^{\frac{a\ln^{7}d}{t}}$
$\leq\exp\left\{\frac{a}{\ln d}\right\}.$
Altogether, we obtain that $|\mathcal{F}|\leq\binom{t}{\log_{C}(\ln^{5}d)}\exp\left\{\frac{a}{\ln d}\right\}\leq\exp\left\{\frac{2a}{\ln d}\right\}$, as required. ∎
We finish this section by showing that given an obstruction with $u\geq 2^{\frac{d}{10C}}$, whp $U$ lies in a set which is not ‘too disconnected’, and whose size is proportional to that of $U$. First, we require the following fairly simple claim.
Claim 3.6.
Whp there are no obstructions such that one (or more) of the following holds,
  1. (a)
    $\ell_{1}\geq 5u$,
  2. (b)
    $s\geq 6u$,
  3. (c)
    $\ell_{3}\geq\frac{6u}{d^{2}}$ and $u\geq d^{2}$.
Proof.
We first claim that in any of the above cases, we would have that
$\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq 5du.$
Indeed, if $\ell_{1}=|V_{1}|\geq 5u$, since the graph $G$ is $d$-regular, $\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq d\ell_{1}\geq 5du$. Similarly, since by Theorem 1.2, every component $K_{i}$ in $G_{p}[S]$ satisfies $e(K_{i},K_{i}^{C})\geq|K_{i}|(d-2(C-1)\log_{2}d)$, if $s\geq 6u$ then $\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq(1-o(1))ds\geq 5du$. Finally, if $\ell_{3}\geq\frac{6u}{d^{2}}$, then, by Theorem 1.2 and by convexity arguments similar to Lemma 3.4, we have that
$\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})$ $>(\ell_{3}-1)d^{2}(d-2(C-1)\log_{2}d)$
$\geq 6du-d^{3}+2(C-1)\log_{2}d\geq 5du,$
where we used the assumption that $u\geq d^{2}$.
Fix $u$ and recall that $U$ has at most $du$ edges touching it. Note that any edge between two components $K_{i},K_{j}$ for $i,j\in[\ell]$ would contradict the fact that we consider an obstruction. Hence, the probability of an obstruction with $\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq 5du$ is at most
$\binom{n}{u}(1-p)^{5du-du}\leq n^{u}n^{-(1-\epsilon)4u}\leq n^{-\frac{5u}{2}},$
where we used $(1-p)^{d}=n^{-(1-\epsilon)}$, and recalling that choosing $U$ determines the obstruction. A union bound over all possible values of $u$, and the at most $n^{2}$ choices for $s$ and $\ell_{3}$ yields that the probability of an event violating the statement of the claim is at most
$3n^{2}\sum_{u=1}^{n}n^{-\frac{5u}{2}}\leq 3n^{2}\cdot 2n^{-\frac{5}{2}}=o(1).$
We are now ready to describe for some range of $u$ a set $M$ containing $U$ with relatively few components.
Lemma 3.7.
Whp for every obstruction with $u\geq 2^{\frac{d}{10C}}$, there is a $M\subseteq V$ satisfying the following,
  1. 1.
    $U\subseteq M$; and,
  2. 2.
    $|M|\leq 10u$; and,
  3. 3.
    $G[M]$ contains at most $u/2^{\frac{d}{(100C)^{3}}}$ components.
Proof.
Suppose that $U$ is an obstruction with $u=|U|$, $|V_{1}|=\ell_{1}$ and $s=|S|$. By Claim 3.6, we have that whp there is no obstruction with $\ell_{1}\geq 4u$, as well as no obstruction with $s\geq 5u$, and no obstruction with $\ell_{3}\geq 10u/d^{2}$. We may thus assume that $\ell_{1}<4u,s<5u$ and $\ell_{3}<10u/d^{2}$. In particular, $\ell_{1}+\ell_{2}\geq u+1-10u/d^{2}$.
For every component $K_{i}$ in $G_{p}[S\cup V_{1}]$, we have that $E_{G_{p}}(K_{i},K_{i}^{C})\subseteq E(K_{i},U)$. Now, by Theorem 1.2,
$\sum_{K_{i}\in G_{p}[S\cup V_{i}]}e(K_{i},K_{i}^{C})\geq d\ell_{1}+(1+o(1))ds\geq d(\ell_{1}+\ell_{2})\geq d(u+1-10u/d^{2}),$
where in the second inequality we used that $s\geq 3\ell_{2}$. Thus, the probability that there is such an obstruction with $e(U,V_{1}\cup S)\leq\frac{du}{100C^{2}}$ is at most
$\binom{n}{u}(1-p)^{d(u+1-10u/d^{2}-\frac{u}{100C^{2}})}$ $\leq\left(\frac{en}{u}\right)^{u}n^{-\left(1-\frac{1}{95C^{2}}\right)u}\leq\exp\left\{\frac{u}{95C^{2}}\ln n+u-u\ln u\right\}$
$\leq\exp\left\{u\left(\frac{\ln n}{95C^{2}}-\frac{\ln 2\cdot d}{10C}\right)\right\}=o\left(n^{-3}\right),$
where we used $n\leq C^{d}$, $C\geq 2$ and $u\geq 2^{\frac{d}{10C}}$. Hence, by the union bound over the at most $n^{3}$ possible choices for $u$, $s$ and $\ell_{3}$, we have that whp any obstruction $U$ with $u\geq 2^{\frac{d}{10C}}$ has $\ell_{1}<4u$, $s<5u$ and there are at least $\frac{du}{100C^{2}}$ edges in $G[U\cup V_{1}\cup S]$.
Let $M=U\cup V_{1}\cup S$. We have that $U\subseteq M$ and $|M|=u+\ell_{1}+s\leq 10u$, by our assumptions $\ell_{1}<4u$ and $s<5u$. 1.2 implies that for any set $M\subseteq V(G)$
$e_{G}(M)\leq|M|d-|M|(d-(C-1)\log_{2}|M|)=|M|(C-1)\log_{2}|M|.$
If $G[M]$ had at least $u/2^{\frac{d}{(100C)^{3}}}$ components, then by standard convexity arguments, we would have that
$e_{G}(M)$ $\leq|M|(C-1)\log_{2}\left(\frac{|M|}{u/2^{\frac{d}{(100C)^{3}}}}\right)$
$\leq 10u(C-1)\log_{2}\left(10\cdot 2^{\frac{d}{(100C)^{3}}}\right)\leq\frac{du}{(100C)^{2}}<\frac{du}{100C^{2}},$
a contradiction. ∎
3.3 Typically no obstructions
We are now ready to show that whp there are no obstructions with $u\geq 1$ in $G_{p}$. We consider several cases separately: when $|U|$ is small; when $s=b=0$; when $s\neq 0$ and $b=0$; when $0\neq b\leq\frac{n}{2}$; and, when $b>\frac{n}{2}$. We show that for each of these cases, whp there are no such obstructions. Note that all events we consider happen with high probability, and thus their intersection also happens with high probability, thus completing the proof of Theorem 3b.
Obstructions with small $U$.
We denote by $\mathcal{B}_{u}$ the event there exists an obstruction in $G_{p}$ with $|U|=u$.
Lemma 3.8.
We have that
$\bigcup_{1\leq u\leq\frac{n}{d^{C^{3}/p}}}\mathbb{P}\left(\mathcal{B}_{u}\right)=o(1).$
Proof.
Suppose there exists a construction with $|U|=u$ and $u\in\left[\frac{n}{d^{C^{3}/p}}\right]$. By Lemma 3.4, we have that
$\sum_{j=1}^{j=w/2}e(W_{j},W_{j}^{C})+\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})$ $\geq w(d-1)+d(\ell-1)+f(k-\ell+1)$
$\geq w(d-1)+du+f(n-2u-w)\geq du+f(2u),$
where we used that $\ell\geq u+1,k=n-u-w$ and the definition of $f(x)$.
We continue by considering two ranges separately. First, suppose that $u\leq 2^{\frac{d}{10C}}$. We then have that $(C-1)\log_{2}u\leq(C-1)\frac{d}{10C}<\frac{d}{10}$ and thus $f(2u)\geq 2u(d-(C-1)\log_{2}u)>\frac{3du}{2}$. As there are at most $du$ edges touching $U$, and noting that any edge between any two different components in $G_{p}[V\setminus U]$ would rule out the existence of such an obstruction, we have that
$\mathbb{P}\left(\mathcal{B}_{u}\right)\leq\binom{n}{u}(1-p)^{3du/2}\leq n^{-(1-\epsilon)\frac{3u}{2}+u}<n^{-u/3}.$
Therefore, by the union bound,
$\bigcup_{1\leq u\leq 2^{\frac{d}{10C}}}\mathbb{P}\left(\mathcal{B}_{u}\right)\leq\sum_{u=1}^{2^{\frac{d}{10C}}}\frac{1}{n^{u/3}}=o(1).$
We now turn to $2^{\frac{d}{10C}}\leq u\leq\frac{n}{d^{C^{3}/p}}$. By Lemma 3.7, we may assume that there exists $M$ such that $|M|\leq 10u$, $U\subseteq M$ and there are at most $u/2^{\frac{d}{(100C)^{3}}}\eqqcolon r$ components in $G[M]$. We have at most $n^{r}$ ways to choose the sizes of these components. Thus, by Lemma 1.3, there are at most $\binom{10u}{u}n^{2r}(ed)^{10u}$ ways to choose $U$. Since $u\leq\frac{n}{d^{C^{3}/p}}$, we have that
$f(2u)\geq\frac{2u(d-\log_{C}u)}{C}\geq\frac{2uC^{2}\log_{C}d}{p}.$
Thus, the probability of such an obstruction is at most
$\binom{10u}{u}n^{2r}(ed)^{10u}(1-p)^{\frac{2uC^{2}\log_{C}d}{p}}$ $\leq\left(10e\right)^{u}n^{2r}(ed)^{10u}\exp\left\{-2uC^{2}\log_{C}d\right\}$
$\leq\exp\left\{u\left(11+\ln 10+\frac{2\ln n}{2^{d/(100C)^{3}}}+10\ln d-2C^{2}\log_{C}d\right)\right\}$
$\leq\exp\left\{u\left(15+\frac{2\ln n}{2^{d/(100C)^{3}}}-\left(10-\frac{2C^{2}}{\ln C}\right)\ln d\right)\right\}$
$=\exp\left\{-\frac{u\ln d}{2}\right\}\leq d^{-u/2},$
using $\frac{2C^{2}}{\ln C}\geq 11$ for $C\geq 2$ and $\frac{2\ln n}{2^{d/(100C)^{3}}}=o(\ln d)$. A union bound over the less than $n$ possible values of $u$ completes the proof. ∎
We now turn to consider obstructions with large $U$. We consider several cases separately. The proofs in Lemmas 3.9-3.14 follow quite similar calculations, however, each such calculation requires a slightly different approach. As these differences are telling of the different challenges in showing that such typical obstructions do not appear, we left the calculations explicit.
Obstructions with $s=b=0$.
Lemma 3.9.
Whp, there are no obstructions with $s=b=0$.
Proof.
Fix $u$ and $w$, and let $\mathcal{B}_{u,w}$ be the event that there is a minimal obstruction with $u=|U|,w=|W|$ and $s=b=0$. By Lemma 3.3, there are at most $2n^{w/2}(ed)^{w/2}$ such minimal obstructions. In particular, $\ell=|V_{1}|=n-u-w$.
We have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq d\ell+w(d-1).$
As there are at most $du$ edges touching $U$ in $G$, we have by the union bound
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}u\in[1,n/2-1]\\ w\in[0,n-2u-1]\end{subarray}}\mathcal{B}_{u,w}\right)$ $\leq\sum_{u=1}^{n/2-1}\sum_{w=0}^{n-2u-1}2n^{w/2}(ed)^{w/2}(1-p)^{d(\ell+w-u)-Cw}$
$\leq\sum_{u=1}^{n/2-1}\sum_{w=0}^{n-3}2n^{-(1-\epsilon)(\ell-u)-w/3}.$
Recall that since $u$ and $w$ are fixed, $\ell-u$ is fixed as well and ranges from $2$ to $n-2$. Furthermore, note that the sum over $w$ is a geometric sum and is at most twice its value when $w=0$. Thus, we have that
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}u\in[1,n/2-1]\\ w\in[0,n-2u-1]\end{subarray}}\mathcal{B}_{u,w}\right)$ $\leq 4\sum_{j=2}^{n-2}n^{-(1-\epsilon)j}=o(1),$
as required. ∎
Obstructions with $s\neq 0,b=0$.
Lemma 3.10.
Whp, there are no obstructions with $s\neq 0$ and $b=0$.
Proof.
Fix $u,w$ and $s$, and let $\mathcal{B}_{u,w,s}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $0\neq s=|S|$ and $b=0$. We then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{j=w/2}e(W_{j},W_{j}^{C})\geq d\ell_{1}+ds-2Cs\log_{2}d+w(d-1),$
since the components in $S$ are of size at most $d^{2}$, and thus by Theorem 1.2, any $K_{i}$ of size at most $d^{2}$ has that $e(K_{i},K_{i}^{C})\geq|K_{i}|(d-2C\log_{2}d)$.
By Lemma 3.3 there are at most $2n^{w/2+\ell_{2}}(ed)^{w/2+2s}$ choices for such a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $b=0$. As before, we obtain that
$\mathbb{P}\left(\bigcup_{w\in[0,n-u-s-\ell_{1}]}\mathcal{B}_{u,w,s}\right)$ $\leq\sum_{w=0}^{n-u-s-\ell_{1}}2n^{w/2+\ell_{2}}(ed)^{w/2+2s}(1-p)^{d(\ell_{1}+s+w-u)-2Cs\log_{2}d-w}$
$\leq 4n^{\ell_{2}}(ed)^{2s}(1-p)^{d(\ell_{1}+s-u)-2Cs\log_{2}d}.$
Now, $\ell_{1}+\ell_{2}\geq u+1$, and thus $\ell_{1}-u\geq 1-\ell_{2}$. Hence,
$n^{\ell_{2}}(ed)^{2s}(1-p$ $)^{d(\ell_{1}+s-u)-2C\log_{2}ds}\leq n^{\ell_{2}}(ed)^{2s}(1-p)^{d(s-\ell_{2}+1)-2Cs\log_{2}d}$
$\leq n^{-(1-\epsilon)(s-2\ell_{2}+1)}(ed)^{2s}(1-p)^{-2Cs\log_{2}d}$
$\leq\exp\left\{-(1-\epsilon)(s-2\ell_{2}+1)\ln n+5C\left(1+\ln\left(\frac{1}{1-p}\right)\right)s\ln d\right\}.$
Note that the above decreases as $s$ increases, and therefore, using $s\geq 3\ell_{2}$ and $\ell_{2}\geq\frac{s}{d^{2}}$, we have that
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}s\in[3,n-u-\ell_{1}]\\ w\in[0,n-u-s-\ell_{1}]\end{subarray}}\mathcal{B}_{u,w,s}\right)$ $\leq 4\sum_{s=3}^{n-u-\ell_{1}}\exp\left\{-(1-\epsilon)(\ell_{2}+1)\ln n+15C\left(1+\ln\left(\frac{1}{1-p}\right)\right)\ell_{2}\ln d\right\}$
$\leq 4d^{2}\sum_{\ell_{2}=1}^{n}\exp\left\{-(1-2\epsilon)(\ell_{2}+1)\ln n\right\}$
$\leq 5d^{2}\exp\left\{-(1-2\epsilon)2\ln n\right\}=o(1/n),$
where we used the fact that $p$ is bounded away from one. Therefore, by the union bound over the at most $n$ values of $u$, we have that whp there is no obstruction violating the statement of the Lemma. ∎
Obstructions with $0\neq b\leq\frac{n}{2}$.
We first treat this case under the assumption that the edge boundary of $B$ in $G$ is not too small.
Lemma 3.11.
Whp there are no obstructions with $0\neq b$ and $e(B,B^{C})\geq b\ln^{2}d$.
Proof.
Fix $u,w,s,\ell_{2},b$, and $\ell_{3}$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $0\neq b=|B|$. Similarly to before, we then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w.$
Since we assume that $e(B,B^{C})\geq b\ln^{2}d$, we obtain that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+b\ln^{2}d+wd-w.$
Now, by Lemma 3.3, there are at most $2n^{w/2+\ell_{2}+2\ell_{3}}(ed)^{w/2+2s+b}=2n^{\ell-\ell_{1}+\ell_{3}+w/2}(ed)^{2s+b+w/2}$ such minimal obstructions with $u=|U|,w=|W|,s=|S|$ and $b=|B|$. As before, we obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2n^{\ell-\ell_{1}+\ell_{3}}(ed)^{2s+b}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+b\ln^{2}d}\cdot n^{w/2}(ed)^{w/2}(1-p)^{dw-w}$
$\leq 2n^{\ell-\ell_{1}+\ell_{3}}(ed)^{2s+b}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+b\ln^{2}d}\cdot n^{-w/3}$
$\leq 2n^{\ell-\ell_{1}+\ell_{3}}(ed)^{2s+b}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+b\ln^{2}d}$
$\leq 2\exp\left\{(\ell-\ell_{1}+\ell_{3})\ln n+(2s+b)\ln(ed)-pb\ln^{2}d\right\}n^{-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)}$
$\leq 2\exp\left\{(\ell-\ell_{1})\ln n+2s(\ln d+1)-\frac{p}{2}b\ln^{2}d-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)\ln n\right\}$
$\leq 2\exp\left\{(\ell+\ell_{3}-(2-\epsilon)\ell_{1}+(1-\epsilon)u)\ln n-s(1-2\epsilon)\ln n-\frac{p}{2}b\ln^{2}d\right\}.$
Recall that $\ell=\ell_{1}+\ell_{2}+\ell_{3}$ and $\ell_{2}\leq\frac{s}{3}$ and $\ell_{3}\leq\frac{b}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{(\ell+\ell_{3}-(2-\epsilon)(\ell-\ell_{2}-\ell_{3})+(1-\epsilon)u)\ln n-s(1-2\epsilon)\ln n-\frac{p}{2}b\ln^{2}d\right\}$
$\leq$ $2\exp\left\{\left((1-\epsilon)\left(u-\ell\right)+\frac{2s}{3}+\frac{3b}{d^{2}}\right)\ln n-s(1-2\epsilon)\ln n-\frac{p}{2}b\ln^{2}d\right\}$
$\leq$ $2\exp\left\{(1-\epsilon)(u-\ell)\ln n+O\left(\frac{b}{d}\right)-s\left(1-2\epsilon-\frac{2}{3}\right)\ln n-\frac{p}{2}b\ln^{2}d\right\}$
$\leq$ $2\exp\left\{O\left(\frac{b}{d}\right)-\frac{p}{2}b\ln^{2}d\right\},$
where in the last step we used that $u<\ell$. Since $b\geq d^{2}$, by the union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$, we have that the probability of an obstruction violating the statement of the claim is at most
$n^{6}\exp\left\{-\Omega\left(d^{2}\ln^{2}d\right)\right\}=o(1).$
We now turn to obstructions where $B$ has a small edge-boundary.
Lemma 3.12.
Whp there are no obstructions with $0\neq b\leq\frac{n}{2}$ and $e(B,B^{C})\leq b\ln^{2}d$.
Proof.
Note that every set $B\subseteq V$ with $e(B,B^{C})\leq b\ln^{2}d$ is of size at least $\frac{n}{d^{\ln^{2}d}}$ since by 1.2, for $b<\frac{n}{d^{\ln^{2}d}}$,
$e(B,B^{C})\geq\frac{b\log_{C}(n/b)}{C}\geq\frac{b\ln^{2}d\log_{C}d}{C}>b\ln^{2}d$
Fix $u,w,s,\ell_{2},b$, and $\ell_{3}$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $0\neq b=|B|\leq\frac{n}{2}$. Similarly to before, we then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+f(b)+wd-w.$
By the proof of Lemma 3.3, there are at most $n^{\ell_{2}+w/2}(ed)^{2s+w/2}$ ways to choose $W\cup S$. Furthermore, since we assume $e(B,B^{C})\leq b\ln^{2}d$ and $b\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2b}{\ln d}\right\}$ choices for $B$. Therefore, by Lemma 3.2, there are at most $2n^{\ell_{2}+w/2}(ed)^{2s+w/2}\exp\left\{\frac{2b}{\ln d}\right\}$ such minimal obstructions. We obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2b}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(b)}\cdot n^{w/2}(ed)^{w/2}(1-p)^{dw-w}$
$\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2b}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(b)}$
$\leq 2\exp\left\{\ell_{2}\ln n+2s\ln(ed)+\frac{2b}{\ln d}-pf(b)\right\}n^{-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)}$
$\leq 2\exp\left\{\ell_{2}\ln n+2s(\ln d+1)+\frac{2b}{\ln d}-pf(b)-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)\ln n\right\}$
$\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell_{1}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2b}{\ln d}-pf(b)\right\}.$
Recall that $\ell=\ell_{1}+\ell_{2}+\ell_{3}$ and $\ell_{2}\leq\frac{s}{3}$ and $\ell_{3}\leq\frac{b}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell-\ell_{2}-\ell_{3}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2b}{\ln d}-pf(b)\right\}$
$\leq 2\exp\left\{\left(\epsilon\frac{s}{3}+(1-\epsilon)\frac{b}{d^{2}}-(1-\epsilon)(\ell-u)\right)\ln n-s(1-2\epsilon)\ln n+\frac{2b}{\ln d}-pf(b)\right\}$
$\leq 2\exp\left\{\frac{3b}{\ln d}-pf(b)\right\},$
where in the last step we used that $u<\ell$. Since we assume that $b\leq\frac{n}{2}$ and $n\leq C^{d}$, we have that $\log_{C}b\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f(b)\geq\frac{b}{C^{2}}$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{b\left(\frac{3}{\ln d}-\frac{p}{C^{2}}\right)\right\}\leq\exp\left\{-\frac{d^{2}}{2C^{2}}\right\}.$
union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},n$ and $\ell_{3}$ completes the proof. ∎
Obstructions with $b>\frac{n}{2}$.
Note that here, by Lemma 3.8, we may assume that $u\geq\frac{n}{d^{C^{3}/p}}$, and thus $|B^{C}|\geq\frac{n}{d^{C^{3}/p}}$. We begin by assuming that $B^{C}$ has a large edge-boundary.
Lemma 3.13.
Whp there are no obstructions with $b>\frac{n}{2}$ and $e(B^{C},B)\geq|B^{C}|\ln^{2}d$.
Proof.
Fix $u,\ell_{1},w,s$ and $b$. We then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w.$
By Claim 3.6, we may assume that $\ell_{3}<\frac{6u}{d^{2}}$, and since $u+1\leq\ell_{1}+\ell_{2}+\ell_{3}$, we have that
$\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w\geq du-\frac{6u}{d}+|B^{C}|\ln^{2}d\geq du+\frac{u\ln^{2}d}{2}.$
We have at most $\binom{n}{u}$ ways to choose the obstruction. Thus, the probability of having an obstruction with such $u,\ell_{1},w,s,b$ is at most
$\binom{n}{u}(1-p)^{u\ln^{2}d/2}$ $\leq\left(\frac{en}{u}\right)^{u}\exp\left\{-\frac{pu\ln^{2}d}{2}\right\}$
$\leq\exp\left\{u\left(1+\ln n-\ln u-\frac{p\ln^{2}d}{2}\right)\right\}$
$\leq\exp\left\{u\left(1+\frac{C^{3}}{p}\ln d-\frac{p\ln^{2}d}{2}\right)\right\}\leq\exp\left\{-\frac{pu\ln^{2}d}{3}\right\},$
where in the penultimate inequality we used $u\geq\frac{n}{d^{C^{3}/p}}$. Union bound over the at most $n^{5}$ choices of $u,\ell_{1},w,s$ and $b$ completes the proof. ∎
We now turn the case where $B^{C}$ has a small edge-boundary.
Lemma 3.14.
Whp there are no obstructions with $b>\frac{n}{2}$ and $e(B^{C},B)\leq|B^{C}|\ln^{2}d$.
Proof.
We follow a similar argument to that in Lemma 3.12. Fix $u,w,s$, and $b$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $\frac{n}{2}\neq b=|B|$. Similarly to before, we then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w.$
By 1.2 $e(B,B^{C})\geq f(b)=f(n-b)$ and we obtain
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+f(n-b)+wd-w.$
By the proof of Lemma 3.3, there are at most $n^{\ell_{2}+w/2}(ed)^{2s+w/2}$ ways to choose $W\cup S$. Furthermore, since we assume $e(B,B^{C})\leq|B^{C}|\ln^{2}d$ and $|B^{C}|\geq u\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2|B^{C}|}{\ln d}\right\}=\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ choices for $B^{C}$, and hence for $B$. Therefore, by Lemma 3.2, there are at most $2n^{\ell_{2}+w/2}(ed)^{2s+w/2}\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ such minimal obstructions. We thus obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(n-b)}\cdot n^{w/2}(ed)^{w/2}(1-p)^{dw-w}$
$\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(n-b)}$
$\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell_{1}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2(n-b)}{\ln d}-pf(n-b)\right\}.$
Recall that $\ell=\ell_{1}+\ell_{2}+\ell_{3}$ and $\ell_{2}\leq\frac{s}{3}$ and by Claim 3.6 $\ell_{3}\leq\frac{6u}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell-\ell_{2}-\ell_{3}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2(n-b)}{\ln d}-pf(n-b)\right\}$
$\leq 2\exp\left\{\left(\frac{6u}{d^{2}}+\epsilon\frac{s}{3}-s(1-2\epsilon)-(1-\epsilon)(\ell-u)\right)\ln n+\frac{2(n-b)}{\ln d}-pf(n-b)\right\}$
$\leq 2\exp\left\{\frac{3(n-b)}{\ln d}-pf(n-b)\right\},$
where in the last step we used that $u<\ell$. Since $b\geq\frac{n}{2}$, we have that $\log_{C}(n-b)\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f(n-b)\geq\frac{n-b}{C^{2}}$. Moreover, $n-b\geq u$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{u\left(\frac{3}{\ln d}-\frac{1}{C^{2}}\right)\right\}\leq\exp\left\{-\frac{u}{2C^{2}}\right\}.$
A union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
4 Discussion
We have extended the classical result concerning hitting times of minimum degree one, connectivity, and the existence of a perfect matching to random subgraphs of regular Cartesian product graphs. In particular, this includes a simplified self-contained version of the connectivity result for bond percolation on the hypercube. Let us mention that, independently, Collares, Doolittle, and Erde use a similar approach – that is, sprinkling with probabilities $p_{1}$ and $p_{2}$ – to show a connectivity result for bond percolation on the permutahedron [9]. There, however, similarly to the approach of [1, 6, 20], one utilises that in $G_{p_{1}}$, large components are relatively well-spread, that is, typically every vertex in $G$ is quite close (in $G$) to a large component of $G_{p_{1}}$. In this paper, we neither require nor utilise such a ‘density’ statement, and instead use the fact that the isolated vertices are ‘sparsely spread’.
We note that Lemma 3.2 does not require anything from the host graph $G$, and Lemma 3.7 only utilises some of the isoperimetric profile of $G$, yet does not utilise the product structure of $G$. Hence, these two results could be of independent interest, in particular in questions concerning the typical existence of a perfect matching under percolation in a more general setting. Still, in this proof, the bound on the number of sets with ‘bad expansion’ from Lemma 3.5 exploits the coordinate structure present in product graphs.
Many other random graph models are known to have typically the same hitting times for minimum degree one, connectivity, and the existence of a perfect matching (see, for example, [15] and the references therein). It is thus natural to ask what are the minimal requirements on $G$ for this phenomenon to hold. As a step towards this, we propose the following question, considering regular graphs with high-degree.
Question 4.1.
Let $G$ be a $d$-regular graph on $n$ vertices, with $d=\omega(1)$ and $n$ divisible by two. What minimal requirements are needed on $G$, such that in the graph process on $G$, the hitting times for minimum degree one, connectivity, and the existence of a perfect matching are the same?
Utilising the Tutte-Berge formula we aimed for a (nearly-)perfect matching, that is, a matching missing none or only one vertex (in the case when $n$ is odd) of the graph. A natural extension is to look for the threshold to have a matching that covers all but a small fraction of the vertices.
Question 4.2.
Let $G=\square_{i=1}^{t}H_{i}$ for $H_{i}$ connected, regular and of bounded size. What is the threshold $p^{*}$ such that for all $p\geq p^{*}$, whp, the giant component of $G_{p}$ contains a (nearly-)perfect matching?
Acknowledgements
The authors would like to thank Joshua Erde, Mihyun Kang and Michael Krivelevich for their guidance, advice, and fruitful discussions. The second author was supported in part by the Austrian Science Fund (FWF) [10.55776/W1230].