### v3

# Perfect Matching in Product Graphs and in their Random Subgraphs

Abstract
For $t\in\mathbb{N}$ and every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph, with $1<|V(H_{i})|\leq C$ for some integer $C\geq 2$. Let $G=\square_{i=1}^{t}H_{i}$ be the Cartesian product of $H_{1},\ldots,H_{t}$. We show that if $t\geq 5C$ then $G$ contains a (nearly-)perfect matching.
Then, considering the random graph process on $G$, we generalise the result of Bollobás on the binary hypercube $Q^{t}$, showing that with high probability, the hitting times for minimum degree one, connectivity, and the existence of a (nearly-)perfect matching in the random graph process on $G$ are the same. As a byproduct, we develop several tools which may be of independent interest in a more general setting when one seeks to establish the typical existence of a perfect matching under percolation.
1 Introduction
1.1 Background and main results
Given two graphs $H_{1}=(V_{1},E_{1})$ and $H_{2}=(V_{2},E_{2})$, the Cartesian product $H=H_{1}\square H_{2}$ is the graph whose vertex set is $V_{1}\times V_{2}$, and for $u_{1},v_{1}\in V_{1}$ and $u_{2},v_{2}\in V_{2}$, we have that $\{(u_{1},u_{2}),(v_{1},v_{2})\}\in E(H)$ either if $u_{1}=v_{1}$ and $\{u_{2},v_{2}\}\in E(H_{2})$ or if $u_{2}=v_{2}$ and $\{u_{1},v_{1}\}\in E(H_{1})$. More generally, given $t$ graphs, $H_{i},\ldots,H_{t}$, their Cartesian product $G=\square_{i=1}^{t}H_{i}$ is the graph with the vertex set
$V\coloneqq\left\{v=(v_{1},\ldots,v_{t})\colon v_{i}\in V(H_{i})\text{ for all }i\in[t]\right\},$
and the edge set
[…]
We call $H_{1},H_{2},\ldots,H_{t}$ the base graphs of $G$. Cartesian product graphs arise naturally in many contexts and have received much attention in combinatorics, probability, and computer science. Many classical graphs, which have been extensively studied, are in fact Cartesian product graphs: the $t$-dimensional torus is the Cartesian product of $t$ copies of the cycle $C_{k}$, the $t$-dimensional grid is the Cartesian product of $t$ copies of the path $P_{k}$, and the binary $t$-dimensional hypercube $Q^{t}$ is the Cartesian product of $t$ copies of a single edge $K_{2}$. We refer the reader to [16] for a systematic coverage of Cartesian product graphs, and related product structures on graphs. Throughout the paper, whenever we write product graphs we consider the Cartesian product as defined above.
In this paper, we study perfect matchings in product graphs. One can convince oneself that if one of the base graphs has a perfect matching, then the product graph $G$ has a perfect matching as well. Indeed, Kotzig [19] showed something even stronger: let $G=\square_{i=1}^{t}H_{i}$ be such that each $H_{i}$ is regular, then if at least one of the base graphs has a $1$-factorisation (recall that a $1$-factorisation is a decomposition of the edges of the graph into disjoint perfect matchings), or at least two base graphs have perfect matchings, then $G$ has a $1$-factorisation; he furthermore demonstrated that while these conditions are sufficient, they are not necessary. Indeed, when none of the base graphs have a perfect matching, it is not clear whether their product graph will have a perfect matching (or nearly-perfect matching, that is, a matching covering all but one vertex when the order of the graph is odd). However, note that given two graphs $H_{1}$ and $H_{2}$, the proportion of vertices covered by a largest matching in $H_{1}\square H_{2}$ is at least as large as this proportion in any of the base graphs. Thus, taking the product of many graphs, one can hope to improve this proportion. Indeed, our first result gives a simple sufficient condition for the existence of a perfect matching in product graphs in this spirit.
Theorem 1.
Let $C>1$ be a constant, and let $t\geq 5C$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$ and suppose that $|V(G)|$ is even. Then $G$ has a perfect matching.
Observe that since each $H_{i}$ is a $d_{i}$-regular connected graph, $G=\square_{i=1}^{t}H_{i}$ is a connected $d$-regular graph with $d=\sum_{i=1}^{t}d_{i}$, and in particular $d=\Theta_{C}(\ln|V(G)|)$. Theorem 1 shows that if the dimension of the product graph is sufficiently large with respect to the maximum size of the base graphs, then the product graph has a perfect matching — regardless of whether the base graphs themselves contain any perfect matchings. Furthermore, we note that the same proof gives that if $|V(G)|$ is odd, then $G$ has a nearly-perfect matching. Finally, the following remark shows that some regularity assumption on the base graphs is necessary.
Remark 1.1.
Note that Theorem 1 is not true if we drop the assumption that the base graphs are regular. Indeed, let $H=K_{1,s}$ be a star and consider the product of $t$ stars $G=\square_{i=1}^{t}H$. Since $K_{1,s}$ is bipartite, $G$ is bipartite for any $t$. However, for any $t\geq 1$, this bipartition is never balanced. In particular, the graph $G$ does not contain a (nearly-)perfect matching. More precisely, denote by $\mathcal{O},\mathcal{E}$ the bipartition sides. Then
$|\mathcal{O}(G)|=\sum_{\begin{subarray}{c}k=0\\ k\text{ even}\end{subarray}}^{t}\binom{t}{k}s^{k}\qquad\text{and}\qquad|\mathcal{E}(G)|=\sum_{\begin{subarray}{c}k=1\\ k\text{ odd}\end{subarray}}^{t}\binom{t}{k}s^{k}.$
The number of vertices in $G$ is $(s+1)^{t}$ and the difference between the bipartition sides is $|\mathcal{O}(G)|-|\mathcal{E}(G)|=(1-s)^{t}$. This is strictly larger than one for any $t\geq 1$ if $s\geq 3$.
As mentioned above, a well-studied product graph is the binary hypercube, $Q^{t}$, in particular in terms of bond percolation on it. Given a graph $G$ and probability $p$, we form the percolated random subgraph $G_{p}\subseteq G$ by including every edge of $G$ independently with probability $p$ (note that $G(n,p)$ is then $(K_{n})_{p}$). The study of $Q^{t}_{p}$ was initiated by Sapoženko [22] and by Burtin [7], who showed that the (sharp) threshold for connectivity is $p^{*}=\frac{1}{2}$: when $p<\frac{1}{2}$ is a constant, with high probability (whp), that is, with probability tending to one as $t$ tends to infinity, $Q^{t}_{p}$ is disconnected, whereas for $p>\frac{1}{2}$ constant, whp $Q^{t}_{p}$ is connected. Erdős and Spencer [14] conjectured that $Q^{t}_{p}$ undergoes a phase transition with respect to its component structure, that is, the typical emergence of a giant component (a connected component containing a linear fraction of the vertices) around $p=\frac{1}{t}$, similar to that of $G(n,p)$ around $p=\frac{1}{n}$. This conjecture was confirmed by Ajtai, Komlós, and Szemerédi [1], with subsequent work by Bollobás, Kohayakawa, and Łuczak [5]. We refer the reader to [20] for a modern short proof of this result.
In recent years, there has been an effort to generalise these results to a wider family of product graphs. Lichev [21] gave sufficient conditions, in terms of the base graphs, for the typical emergence of a giant component in bond percolation on high-dimensional product graphs. Diskin, Erde, Kang, and Krivelevich [10] improved upon this, giving sufficient and tight conditions for the typical emergence of a giant component. Furthermore, they showed that assuming the base graphs are regular, one can give a rather precise description, similar to that in $G(n,p)$, both of the typical component structure [11], and of the asymptotic combinatorial properties of the giant component [9].
Returning to the dense regime, that is, when $p$ is constant, connectivity and the existence of a perfect matching in $Q^{t}_{p}$ have been studied in detail. In particular, Bollobás [4] obtained a hitting time result for the random graph process on $Q^{t}$. Given a graph $\Gamma$, the random graph process on $\Gamma$ is defined as a random sequence of nested graphs $\Gamma(0)\subseteq\ldots\subseteq\Gamma(|E(\Gamma)|)$ together with an ordering $\sigma$ on $E(\Gamma)$, chosen uniformly at random from among all $|E(\Gamma)|!$ such orderings. We set $\Gamma(0)$ to be the empty graph on $V(\Gamma)$. Given $\Gamma(i)$, with $0\leq i<|E(\Gamma)|$, we form $\Gamma(i+1)$ by adding the $(i+1)$-th edge, according to the ordering $\sigma$, to $\Gamma(i)$. The hitting time of a monotone increasing, non-empty graph property $\mathcal{P}$, is the random variable equal to the index $\tau$ for which $\Gamma(\tau)\in\mathcal{P}$, but $\Gamma(\tau-1)\notin\mathcal{P}$. Note that having minimum degree one, connectivity, and the existence of a perfect matching are all monotone increasing properties. Furthermore, observe that for a graph $\Gamma$ to be connected or to contain a perfect matching, the minimum degree of $\Gamma$ has to be at least one. A classical result of Erdős and Rényi [13], and of Bollobás and Thomason [6], is that in the random process on $K_{n}$, whp the hitting time for minimum degree one, connectivity, and the existence of a perfect matching are the same. In 1990, Bollobás [4] showed that the same phenomenon holds in the random graph process on $Q^{t}$.
Theorem 1.2.
Consider the random graph process on $Q^{t}$. Let $\tau_{1}$ be the hitting time of minimum degree one, let $\tau_{2}$ be the hitting time for connectivity, and let $\tau_{3}$ be the hitting time for the existence of a perfect matching. Then, whp, $\tau_{1}=\tau_{2}=\tau_{3}$.
Subsequent work by Joos determined the threshold for connectivity for Cartesian powers of graphs [17, 18], that is, for $G=\square_{i=1}^{t}H$ where all the base graphs $H_{i}=H$ are the same.
Our main result generalises the result of Bollobás [4] to a wider family of product graphs.
Theorem 2.
Let $C>1$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph, with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$, and suppose that $|V(G)|$ is even. Consider the random graph process on $G$. Let $\tau_{1}$ be the hitting time of minimum degree one, let $\tau_{2}$ be the hitting time for connectivity, and let $\tau_{3}$ be the hitting time for the existence of a perfect matching. Then, whp, $\tau_{1}=\tau_{2}=\tau_{3}$.
In fact, we prove the following description of $G_{p}$, from which we derive Theorem 2.
Theorem 3.
Let $C>1$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$, and suppose that $n\coloneqq|V(G)|$ is even. Let $d\coloneqq d(G)=\sum_{i=1}^{t}d_{i}$. Let $\omega=\omega(d)$ be a function tending arbitrarily slowly to infinity, and let $p$ be such that $(1-p)^{d}=\omega/n$. Then, whp, the following holds in $G_{p}$.
  1. (a)
    There exists a unique giant component, spanning all but $o(n)$ of the vertices. All the other components of $G_{p}$, if there are any, are isolated vertices. Furthermore, every two isolated vertices in $G_{p}$ are at distance at least two in $G$.
  2. (b)
    The giant component of $G_{p}$ has a (nearly-)perfect matching.
A few comments are in place. Note that the threshold probability $p^{*}$, at which the expected number of isolated vertices is zero, is $p^{*}=1-\left(\frac{1}{n}\right)^{1/d}$. Noting that $t\leq d\leq(C-1)t$, we have that $d=\Theta_{C-1}(\ln n)$. Thus, $\left(\frac{1}{n}\right)^{1/d}$ is bounded away from zero, and bounded away from one, and in turn, $p^{*}$ is bounded away both from zero and from one. Further, if $p$ is such that $(1-p)^{d}=\omega/n$ for $\omega$ tending to infinity arbitrarily slowly, then $p$ is bounded away both from zero and from one.
Note, that for our choice of $p$ such that $(1-p)^{d}=\omega/n$, whp, there are still isolated vertices. Furthermore, let us remark that Theorem 3a implies that whp adding any edge to $G_{p}$, which is not contained in the giant component, must connect an isolated vertex to the giant component, and indeed this suffices to show that the hitting times of minimum degree one and connectivity are whp the same. Theorem 3b then shows that whp the only obstacle to a perfect matching in $G_{p}$ itself is the existence of isolated vertices. Thus, by standard coupling between the random graph process and the bond percolation model, Theorem 2 indeed follows from Theorem 3.
Let us further remark that the proof of Theorem 3a does not rely on the product structure of the graph, and only uses the assumptions that the graph is $d$-regular, has optimal edge-expansion for sets of size polynomial in $d$, and has some mild edge-expansion for larger sets. Only in the proof of Theorem 3b we use the product structure explicitly. Moreover, since the hypercube is bipartite, in order to find a perfect matching in $Q^{t}_{p}$, it suffices to verify Hall’s condition. In our setting, however, the graph $G$ is not necessarily bipartite, which means that in order to show the typical existence of a perfect matching, one needs to consider the Tutte-Berge formula. This requires a much more delicate treatment, which is detailed in the proof outline in the subsequent section. This further requires us to develop tools which hold for wider family of graphs (see Lemmas 3.2, 3.3, and 3.4). These could be of independent interest when treating perfect matchings in the setting of percolation on regular graphs.
The paper is structured as follows. In Section 1.2 we give an outline of the proofs. In Section 2 we prove Theorem 3a, and in Section 3 we prove Theorem 1 and Theorem 3b. Finally, in Section 4 we conclude with a short discussion of the results and possible avenues for future research.
1.2 Proofs outline
Notation and definitions.
Throughout the paper we let $C>1$ be an integer, $(H_{i})_{i=1}^{t}$ be a sequence of $d_{i}$-regular and connected graphs with $1<|V(H_{i})|\leq C$, and let $G=\square_{i=1}^{t}H_{i}$ be their product. We let $G=(V,E)$. We call $t$ the dimension of $G$, and given $u=(u_{1},\ldots,u_{t})\in V$ we call $u_{i}$ the $i$-th coordinate of $u$. Furthermore, we denote by $n:=|V|$ the order of $G$, and by $d\coloneqq d(G)$ the degree of $G$, noting once again that $d=\sum_{i=1}^{t}d_{i}=\Theta_{C}(\ln n)$.
Given a graph $\Gamma$ and subsets $S_{1},S_{2}\subseteq V(\Gamma)$ with $S_{1}\cap S_{2}=\varnothing$, we denote by $E_{\Gamma}(S_{1})$ the set of edges in $\Gamma$ whose both endpoints are in $S_{1}$, and by $E_{\Gamma}(S_{1},S_{2})$ the edges in $\Gamma$ with one endpoint in $S_{1}$ and the other endpoint in $S_{2}$. We set $e_{\Gamma}(S_{1})\coloneqq|E_{\Gamma}(S_{1})|$ and $e_{\Gamma}(S_{1},S_{2})\coloneqq|E_{\Gamma}(S_{1},S_{2})|$. If the choice of $\Gamma$ is clear, we may omit the subscript. Moreover, given $S\subseteq V(\Gamma)$, we define $S^{C}$ to be $V(\Gamma)\setminus S$.
External results.
We make extensive use of the following generalisations of Harper’s inequality to regular high-dimensional product graphs.
Theorem 1.3 (Theorems 1 and 2 in [9] and Section 3.4 in [12]).
For every $S\subseteq V$ with $1\leq|S|\leq n$,
$e(S,S^{C})$ $\geq\frac{|S|\cdot e\cdot\ln\left(\frac{n}{|S|}\right)}{C}\qquad\text{ and }$
$e(S,S^{C})$ $\geq|S|(d-(C-1)\log_{C}|S|).$
We also utilise the following bound on the number of trees on $k$ vertices in a $d$-regular graph, that are rooted at a fixed vertex.
Lemma 1.4 (Lemma 2 in [3]).
Let $\Gamma$ be a $d$-regular graph, let $k$ be a positive integer and let $v\in V(\Gamma)$. Denote by $t_{k}(v)$ the number of trees on $k$ vertices rooted at $v$ in $\Gamma$. Then $t_{k}(v)\leq(ed)^{k-1}$.
Proof outline of Theorem 3a.
The proof follows from a double-exposure argument, similar in spirit to the classical argument of Ajtai, Komlós, and Szemerédi [1]. Let $p$ be as in the statement of Theorem 3. Let $p_{2}\coloneqq\frac{1}{d^{2}}$ and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note that $G_{p_{1}}\cup G_{p_{2}}\sim G_{p}$, and $p_{1}\approx p-\frac{1}{d^{2}}$. We first show, using a first-moment calculation, that whp every two vertices $u,v$ which are isolated in $G_{p_{1}}$ are at distance at least two in $G$ (Lemma 2.1). Then, utilising Theorem 1.3 and Lemma 1.4, we show that whp there are no components in $G_{p_{1}}$ whose order is in $\left[2,d^{20}\right]$ (Lemma 2.2). We then turn to show that typically components whose order is at least $d^{20}$ merge after sprinkling with $p_{2}$ — this is fairly standard, and follows the same approach as in [1]. Noting that any vertex which was an isolated vertex in $G_{p_{1}}$, is whp either an isolated vertex after sprinkling with $p_{2}$ or merged into a component which was of size at least $d^{20}$ in $G_{p_{1}}$, we obtain that whp $G_{p}$ has a unique large connected component, and all the other vertices are isolated vertices. Finally, from Markov’s inequality it follows that the total volume of isolated vertices in $G_{p}$ is $o(n)$.
Proof outline of Theorem 3b.
Recall that we seek to show that the giant component of $G_{p}$ has a (nearly-)perfect matching. We assume divisibility by $2$ whenever necessary, noting that minor modifications will allow us to obtain a nearly-perfect matching otherwise. We know from Theorem 3a that whp every vertex of $G$ which is not in the giant component is an isolated vertex in $G_{p}$. Thus, it suffices to show that whp $G_{p}$ has a perfect matching if there are typically no isolated vertices in $G_{p}$. To that end, we will utilise the Tutte-Berge formula [2], which implies that the number of unmatched vertices in a maximum matching in a graph $\Gamma$ is equal to
$\max_{U\subseteq V(\Gamma)}\left(\mathrm{odd}\left(\Gamma[V(\Gamma)\setminus U]\right)-|U|\right),$
where $\mathrm{odd}\left(\Gamma[V(\Gamma)\setminus U]\right)$ is the number of connected components with an odd number of vertices in $\Gamma[V(\Gamma)\setminus U]$. In particular, a graph $\Gamma$ has a perfect matching if and only if for every subset $U\subseteq V(\Gamma)$, the subgraph $\Gamma[V(\Gamma)\setminus U]$ has at most $|U|$ connected components with an odd number of vertices. Note here that if $\Gamma$ has isolated vertices, then choosing $U=\varnothing$ witnesses that there is no perfect matching in $\Gamma$.
We thus define Tutte-like obstructions — we say that $U\subseteq V$ is an obstruction (for $G_{p}$) if $|U|\geq 1$, and the number of components of size different than two in $G_{p}[V\setminus U]$ is at least $|U|+1$. We say that the size of the obstruction is $|U|$. Note that if $G_{p}$ has no such obstructions, then the only possible obstructions to a perfect matching are isolated vertices, and whp the giant component has a perfect matching. Furthermore, observe that this definition does not capture all the properties coming from the Tutte-Berge formula – indeed, we treat components with an even number of vertices, except two, the same as components with an odd number of vertices – but this crude, yet much-simplified, outlook will (almost) suffice for our needs.
We say that $U$ is a minimal obstruction, if it is an obstruction with the smallest size of $U$. Note that if a graph has an obstruction, then it must also have a minimal obstruction.
We will use the convention that $u\coloneqq|U|$. In order to obtain a finer description of the components in $G_{p}[V\setminus U]$ we denote the components of size two in $G_{p}[V\setminus U]$ by $W_{1},\ldots,W_{w/2}$, such that their total volume is $w$. Let us partition the vertices in the components that are not of size two into three sets, according to the size of the component in $G_{p}[V\setminus U]$ they belong to. Let $\ell_{1}$ be the number of components of size one in $G_{p}[V\setminus U]$, let $\ell_{2}$ be the number of components whose size is in $3,\frac{n}{d^{C^{3}/p}}]$ in $G_{p}[V\setminus U]$, and let $\ell_{3}=\ell-\ell_{1}-\ell_{2}$ be the number of components in $G_{p}[V\setminus U]$ that contain more than $\frac{n}{d^{C^{3}/p}}$ vertices. We further denote the set of vertices in components of $G_{p}[V\setminus U]$ whose size is one by $V_{1}$, those whose size is in $3,\frac{n}{d^{C^{3}/p}}]$ by $S$, and those whose size is larger than $\frac{n}{d^{C^{3}/p}}$ by $B$. We write $s:=|S|$ and $b:=|B|$, and note that $\ell_{1}=|V_{1}|$.
We remark that the choice of the values for defining $S$ and $B$ is not (completely) arbitrary, and we will see these come into play throughout the proof.
Figure 1: Illustration of an obstruction, with the sets $U$, $V_{1}$, $W$, $S$ and $B$. Note that the only edges in $G_{p}$, which are not induced by $U$ or by components from $W,S$ or $B$, are in $E(U,V_{1}\cup W\cup S\cup B)$.
Using 1.3 one can deduce a lower bound on the number of edges leaving components in $G_{p}[V\setminus U]$. Note that for $U$ to be an obstruction, for any component $K$ of $G_{p}[V\setminus U]$ the edges $E_{G_{p}}(K,K^{C})\subseteq E(K,U)$. Since $G$ is $d$-regular, there are at most $du$ edges touching $U$. So in order to have an obstruction in $G_{p}$, many of the edges leaving components in $G_{p}[V\setminus U]$ cannot be present in $G_{p}$.
We will then treat minimal obstructions differently, depending on their size (that is, on the size of $u$). For $u\leq d$, we will have a small caveat. We call an obstruction $U$ with $\ell=u+1,\ell_{1}=u$ and $\ell_{2}+\ell_{3}=1$ a trivial obstruction. Note that any trivial obstruction is not a real obstruction to a perfect matching in the Tutte-Berge sense as the single component in $S\cup B$ is actually of even size ($n-w-\ell_{1}-u$ is even since $n,w$ are even and $\ell_{1}=u$). Thus, we may assume that any obstruction we consider is not a trivial obstruction. Using this observation, we will be able to show that whp there are no minimal non-trivial obstructions when $u\leq d$ (Lemma 3.10).
Turning to obstructions of more substantial size, we note that an obstruction is uniquely determined by the choice of $U$, and that $U$ is not necessarily connected, and thus the number of choices could be $\binom{n}{u}$. As is often the case, a key ingredient will be to efficiently enumerate the number of possible obstructions, and $\binom{n}{u}$ will often be an inefficient bound. We thus prove several claims, which could be of independent interest in arguing about perfect matchings under percolation for general graphs. In Lemma 3.3, we show that a minimal obstruction of size $u$ is determined by the choice of $W\cup S\cup B$. We stress that this result holds for any graph $G$, without requirements on its degree or structure. As these sets contain connected components, by Lemma 1.4 there are at most (roughly) $n^{w/2+\ell_{2}+\ell_{3}}(ed)^{s+w+b}$ choices for a minimal obstruction of size $u$ (see Lemma 3.4 for a precise statement). Since $d=\Theta_{C}(\ln n)$, for ‘large’ $u$ this will be a much more efficient bound than $\binom{n}{u}$. Furthermore, if $U$ is a minimal obstruction, then every vertex $v\in U$ has neighbours in at least three components of $G_{p}[V\setminus U]$ (see Lemma 3.2). For certain values of $u,\ell_{1},\ell_{2},\ell_{3}$, this implies that typically an obstruction $U$ is contained in a well-connected set of the same order (see Lemma 3.14 for a precise statement). This allows for an even more accurate enumeration. This more effecient counting approach will allow us to treat obstructions up to size $u=n/poly(d)$.
For larger values of $u$, the aforementioned probabilistic bound no longer suffices. Thus, the final ingredient, similar in spirit to the approach of Bollobás in [4], will be Lemma 3.5, which allows to efficiently bound the number of choices for sets $B$ that have ‘bad expansion’, that is, the number of edges in $E(B,B^{C})$ is small. We note that this is the only place where we use the structure of $G$ as a product graph. Indeed, sets $B$ with bad expansion leave a large ‘fingerprint’ on some coordinates of the product, in particular, if $B$ intersects non-trivially with a projection $G(I)$ of $G$ onto some coordinates $I\subseteq[t]$, then $G(I)$ spans many edges of $E(B,B^{C})$.
Finally, let us briefly comment on the difference between the proof here and the proof for the hypercube, given in [4]. Since the hypercube is a bipartite graph, it suffices there to apply Hall’s theorem, and in particular, one does not need to consider the entirety of the giant component, but only sets of size at most $\frac{n}{2}$ which expand badly. Hence, Harper’s isoperimetric inequality for the hypercube, together with a fingerprint argument given by Bollobás [4], allows one to show that whp there are no obstructions (bar isolated vertices) to a perfect matching. Here, since $G$ is not necessarily bipartite, one needs to consider Tutte-like obstructions, and in particular sets encompassing the entirety of the giant component. This, in turn, requires a much more delicate treatment, and in particular, to consider the structure of obstructions, and the more efficient enumerations given by Lemmas 3.3 and Lemma 3.4 as well as the fact that for a certain range of parameters an obstruction $U$ is typically contained in a well-connected set of the same order (see Lemma 3.14 for a precise statement). This part of the analysis is the key novelty of this work.
2 Connectivity
Throughout this section, let $\omega=\omega(d)$ be a function tending arbitrarily slowly to infinity and let $p$ be such that $(1-p)^{d}=\omega/n$. We begin by showing that isolated vertices in $G_{p}$ are typically at distance at least two in $G$.
Lemma 2.1.
Whp, every two isolated vertices in $G_{p}$ are at distance at least two.
Proof.
Fix an edge $\{u,v\}\in E$. Since $G$ is $d$-regular, the total number of edges meeting $u$ and $v$ is 2d12d-1. Thus, the probability that $u$ and $v$ are isolated vertices in $G_{p}$ is at most
$(1-p)^{2d-1}\leq(1-p)^{-1}\left(\frac{\omega}{n}\right)^{2},$
where we used that $(1-p)^{d}=\omega/n$. There are $\frac{dn}{2}$ edges to consider. Thus, by the union bound, the probability that two isolated vertices in $G_{p}$ are at distance one in $G$ is at most $\frac{dn}{2}\left(\frac{\omega}{n}\right)^{2}(1-p)^{-1}=o(1)$. ∎
Let us further show the following ‘gap’ statement, which is an almost immediate corollary of Theorem 1.3 and Lemma 1.4. We note that $d^{20}$ in the following lemma is chosen rather arbitrarily, and a much tighter bound can be obtained.
Lemma 2.2.
Whp, there are no connected components $K$ in $G_{p}$ with $|V(K)|\eqqcolon k\in[2,d^{20}]$.
Proof.
Fix $k\in[2,d^{20}]$. Let us estimate the probability that there is a connected component $K$ of order $k$ in $G_{p}$. By Theorem 1.3, we have that $e(K,K^{C})\geq k(d-(C-1)\log_{C}k)\geq\frac{9kd}{10}$, where we used the fact that $k\leq d^{20}$. Let $\mathcal{T}_{k}$ be the set of trees of order $k$, and let $\mathcal{A}_{k}$ be the event that there exists a connected component of order $k$ in $G_{p}$. Thus, by the union bound and by Lemma 1.4,
$\mathbb{P}\left[\mathcal{A}_{k}\right]$ $\leq\sum_{T\in\mathcal{T}_{k}}(1-p)^{e(T,T^{C})}\leq n(ed)^{k-1}(1-p)^{\frac{9kd}{10}}\leq n\left(\frac{\omega}{n}\right)^{9k/10}(ed)^{k-1}\leq n^{1-4k/5}.$
By the union bound over the less than $d^{20}$ possible values of $k$, we have that the probability of an event violating the statement of the lemma is at most
$\sum_{k=2}^{d^{20}}n^{1-4k/5}\leq 2n^{1-8/5}\leq n^{-1/2}=o(1).$
We are now ready to prove the key result of this section, that is, Theorem 3a, whose proof will utilise the classical double-exposure argument of [1].
Proposition 2.3.
Whp, there is a unique giant component in $G_{p}$ whose order is no(n)n-o(n), and all the other components in $G_{p}$ are isolated vertices. Furthermore, every two isolated vertices are at distance at least two from each other.
Proof.
Let $p_{2}=\frac{1}{d^{2}}$, and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note that there exists a function $\omega^{\prime}$ tending to infinity such that $(1-p_{1})^{d}=\omega^{\prime}/n$, and thus we may apply Lemmas 2.1 and 2.2 on $G_{p_{1}}$. Furthermore, observe that $G_{p}$ has the same distribution as $G_{p_{1}}\cup G_{p_{2}}$.
We begin by percolating with probability $p_{1}$. Note that by Lemma 2.2, whp in $G_{p_{1}}$ there are only isolated vertices and components of order at least $d^{20}$. We begin by showing that after sprinkling with $p_{2}$, typically all the components of order at least $d^{20}$ merge. Let $L$ be the set of vertices in components of order at least $d^{20}$ in $G_{p_{1}}$. Let $A\sqcup B=L$ be a partition of $L$ which respects the components of $G_{p_{1}}$, that is, every component is contained either in $A$ or in $B$. We may assume that $|A|\leq|B|$ and let $a\coloneqq|A|$.
Let $A^{\prime}\coloneqq\left(A\cup N_{G}(A)\right)\setminus B$, where $N_{G}(X)$ is the external neighbourhood of $X$ in $G$, and let $B^{\prime}\coloneqq\left(B\cup N_{G}(B)\right)\setminus A^{\prime}$. By Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are at distance at least two from each other in $G$. We continue assuming this holds deterministically. Thus, every isolated vertex in $G_{p_{1}}$ is in the neighbourhood of $A$ or $B$, and we have that $V=A^{\prime}\sqcup B^{\prime}$, that is, $B^{\prime}=V\setminus A^{\prime}$. By Theorem 1.3, we have that $e(A^{\prime},B^{\prime})\geq\frac{a\cdot e\cdot\ln(n/a)}{C}\geq\frac{a\cdot e\cdot\ln 2}{C}$, since $a\leq\frac{n}{2}$. We can extend these edges to $\frac{a}{C}$ paths of length at most $3$ between $A$ and $B$. Every two such paths share at most two edges. Thus, we can trim the set of $\frac{a}{C}$ paths to a set of $\frac{a}{Cd^{2}}$ edge-disjoint paths between $A$ and $B$.
We now sprinkle with probability $p_{2}$. The probability none of these paths are in $G_{p_{2}}$ is thus at most $(1-p_{2}^{3})^{\frac{a}{Cd^{2}}}\leq\exp\left\{-\frac{a}{d^{9}}\right\}$. For a union bound note that there are at most $\binom{n/d^{20}}{a/d^{20}}$ possibilities to choose the partition of $L$. Thus, the probability that there is a partition of $L$ such that there is no path between two components of order at least $d^{20}$ is at most
$\sum_{a=d^{20}}^{\frac{n}{2d^{20}}}\binom{n/d^{20}}{a/d^{20}}\exp\left\{-\frac{a}{d^{9}}\right\}$ $\leq\sum_{a=d^{20}}^{n}(en)^{a/d^{20}}\exp\left\{-\frac{a}{d^{9}}\right\}=o(1),$
where we used $\binom{m}{k}\leq(em)^{k}$ and the fact that $\ln n\leq\ln(C^{d})\leq d^{2}$, as well as that $a\geq d^{20}$. Thus, the probability that there is a component of order at least $d^{20}$ which does not merge in $G_{p_{1}}\cup G_{p_{2}}$ tends to zero. That is, all the components of order at least $d^{20}$ in $G_{p_{1}}$ merge after sprinkling with probability $p_{2}$.
Note that by Lemma 2.1, whp every two isolated vertices in $G_{p_{1}}$ are not connected by an edge of $G$. Hence, adding any edge touching an isolated vertex connects it to a component whose order is at least $d^{20}$ in $G_{p_{1}}$, and these components all merge whp. Hence, whp, there exists a unique connected component in $G_{p_{1}}$ whose order is at least $d^{20}$, and all the other components are isolated vertices, whose distance in $G$ is at least two. Finally, let $Y$ be the random variable counting the number of isolated vertices in $G_{p}$. Then, $\mathbb{E}[Y]=n(1-p)^{d}\leq\omega$. Thus, by Markov’s inequality, whp $Y\leq\omega^{2}=o(n)$, completing the proof. ∎
3 Perfect matching
This section is structured as follows. In Subsection 3.1 we prove Theorem 1. Then, in Subsection 3.2 we prove several estimates on the number of obstructions and related sets. Finally, in Subsection 3.3 we prove Theorem 3b. Throughout the section, we assume divisibility by two whenever is necessary, noting that minor modifications will allow us to argue for the existence of nearly-perfect matching otherwise.
3.1 Proof of Theorem 1
We begin with the proof of Theorem 1. Let us first show that, by Theorem 1.3, for ‘high-enough’ dimension of $G$ every edge-cut of $G$ has at least $d$ edges.
Lemma 3.1.
Suppose that $t\geq 5C$. Then, we have that $G$ is $d$-edge-connected.
Proof.
It suffices to show that $e(S,S^{C})\geq d$ for all $S\subseteq V$ with $1\leq|S|\leq n-1$, and in fact by symmetry, for $S$ of size at most $\frac{n}{2}$. By 1.3,
$e(S,S^{C})\geq|S|(d-(C-1)\log_{C}|S|).$ (1)
Taking a derivative in $|S|$ we obtain, that for all $|S|$ such that $d\geq(C-1)\log_{C}|S|+(C-1)/\ln 2$, the right-hand side of (1) is increasing with $|S|$. Consider $|S|\leq d$. Then,
$d-(C-1)\log_{C}(|S|)\geq d-(C-1)\log_{C}d>2C\,$,
where we used that $d\geq t\geq 5C$ and $C\geq 2$. Hence, $|S|(d-(C-1)\log_{2}|S|)$ is increasing with $|S|$ for $|S|\leq d$, and for all $S$ with $|S|\leq d$ we have that $e(S,S^{C})\geq d$, as required. For $d\leq|S|\leq C^{\frac{d-1}{C-1}}$ we have $d-(C-1)\log_{C}|S|\geq 1$, and thus $e(S,S^{C})$ is at least $d$, as required. Thus, suppose that $C^{\frac{d-1}{C-1}}\leq|S|\leq\frac{n}{2}$. Note that $|S|\leq n/2$, which implies that $\ln(n/|S|)\geq\ln 2$. Hence, by 1.3,
$e(S,S^{C})$ $\geq\frac{|S|\cdot e\cdot\ln(n/|S|)}{C}\geq\frac{|S|}{C}\geq\frac{C^{\frac{d-1}{C-1}}}{C}=C^{\frac{d-1}{C-1}-1}\,.$ (2)
We claim that for our choice of $t$, (2) is at least $d$. Note that the $C^{\frac{d-1}{C-1}}-1-d$ is increasing as $d$ is increasing, and indeed
$C^{\frac{5C-1}{C-1}-1}-5C>C^{3}-5C\geq 0,$
for $C\geq 2$. ∎
The proof of Theorem 1 then follows immediately from Lemma 3.1, since every $d$-regular, (d1)(d-1)-edge-connected graph has a perfect matching (see [2, Chapter 18, Theorem 7]).
3.2 Properties of obstructions
Recall from Section 1.2 the definition of a minimal obstruction, further recall the sets $U,V_{1},W,S$, and $B$, and $\ell,\ell_{1},\ell_{2},\ell_{3}$. In this Section, we obtain several deterministic lemmas, which will allow us to count the number of minimal obstructions more efficiently.
The first lemma is a rather simple (yet, as we will see in Lemma 3.13, quite powerful) observation.
Lemma 3.2.
Let $U$ be a minimal obstruction. Then any vertex $v\in U$ has neighbours in at least three components of $G_{p}[V_{1}\cup S\cup B]$.
Proof.
Suppose towards contradiction that there is a minimal obstruction $U$ of size $u$ with $\ell>u$ components in $G_{p}[V_{1}\cup S\cup B]$ containing a vertex $v$ whose neighbourhood intersects at most two components of $G_{p}[V_{1}\cup S\cup B]$.
Consider $U^{\prime}=U\setminus\{v\}$. Clearly, $|U^{\prime}|=u-1$. Let us denote by $V_{1}^{\prime},S^{\prime},B^{\prime}$ the sets containing (respectively) components of size exactly one, between $3$ and $\frac{n}{d^{C^{3}/p}}$, and larger than $\frac{n}{d^{C^{3}/p}}$ in $G_{p}[V\setminus U^{\prime}]$. Observe that since $v$ is connected to at most two components in $G_{p}[V_{1}\cup S\cup B]$, there are at least $\ell-1>u-1=|U^{\prime}|$ components in $G_{p}[V_{1}^{\prime}\cup S^{\prime}\cup B^{\prime}]$. But then $U^{\prime}$ is an obstruction, contradicting the minimality of $U$. ∎
The following observation will be used in the subsequent lemma, Lemma 3.4, to count obstructions efficiently by choosing $W,S$ and $B$ (instead of $U$).
Lemma 3.3.
There are at most two minimal obstructions, $U$ and $U^{\prime}$, with $|U|=|U^{\prime}|=u\geq 2$ and the same choice of $W\cup S\cup B$.
Proof.
Fix a minimal obstruction $U$ with $u=|U|\geq 2$. We have that $V\setminus(U\cup W\cup S\cup B)=V_{1}$.
Suppose first that there are two sets, $\varnothing\neq A_{1}\subsetneq U$ and $\varnothing\neq A_{2}\subseteq V_{1}$, such that $X\coloneqq(U\setminus A_{1})\cup A_{2}$ is a minimal obstruction as well. By minimality, we have that $|A_{1}|=|A_{2}|$. Since $X$ is also an obstruction and has the same set of components of size strictly larger than one, every $v\in V_{1}\setminus A_{2}$ has that $N(v)\subseteq U\setminus A_{1}$. But then, we have that $U\setminus A_{1}$ is also an obstruction, where $1\leq|U\setminus A_{1}|<u$ — contradicting the minimality of $U$. Indeed, there are $\ell_{1}-|A_{2}|$ components of size $1$ in $G_{p}[V\setminus(U\setminus A_{1})]$, and $\ell_{2}+\ell_{3}$ components of size at least three in $G_{p}[S\cup B]$, thus at least $\ell_{1}+\ell_{2}+\ell_{3}-|A_{2}|=\ell-|A_{2}|\geq u+1-|A_{2}|=|X|+1$ components in total.
Now, suppose towards contradiction that there are three minimal obstructions $U,U^{\prime},U^{\prime\prime}$ with the same choice of $W\cup S\cup B$. Then, by the above, we have $\varnothing\neq U^{\prime},U^{\prime\prime}\subseteq V_{1}$ with $U^{\prime}\neq U^{\prime\prime}$, such that $u=|U^{\prime}|=|U^{\prime\prime}|$. But then, note that there must be some $v\in U^{\prime\prime}\subseteq V_{1}$, with $v\notin U^{\prime}$. Since $N(v)\subseteq U$, $N(v)\subseteq U^{\prime}$, and $U^{\prime}\cap U=\varnothing$, we have that $v$ is an isolated vertex. But then $U^{\prime\prime}\setminus\{v\}$ is also an obstruction of size $u-1\geq 1$ — contradiction. ∎
We will often seek to enumerate the number of minimal obstructions, having fixed $w,\ell_{2},s,\ell_{3},$ and $b$.
Lemma 3.4.
Given $u,w,\ell_{2},s,\ell_{3},$ and $b$, the number of minimal obstructions is at most
$2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$
Proof.
Recall that $W$ is a set of $w/2$ edges. We thus have at most $\binom{nd/2}{w/2}\leq(nd)^{w/2}$ ways to choose $W$. As for $S$, it has $\ell_{2}$ components. Let us denote the sizes of these components by $s_{1},\ldots,s_{\ell_{2}}$, where we have that $3\leq s_{i}\leq\frac{n}{d^{C^{3}/p}}$ for every $i\in[\ell_{2}]$. Thus there are at most $(\frac{n}{d^{C^{5}/p}})^{\ell_{2}}\leq n^{\ell_{2}}$ ways to choose $s_{1},\ldots,s_{\ell_{2}}$. We then have at most $\binom{n}{\ell_{2}}\leq n^{\ell_{2}}$ ways to choose roots for some spanning trees of these components. Thus by Lemma 1.4, the number of ways to choose $S$ is at most $n^{2\ell_{2}}\prod_{i=1}^{\ell_{2}}(ed)^{s_{i}-1}\leq n^{2\ell_{2}}(ed)^{s}$. Finally, $B$ has $\ell_{3}$ components, and there are at most $n^{\ell_{3}}$ ways to choose their sizes. Then, similarly to the above, by Lemma 1.4 there are at most $n^{2\ell_{3}}(ed)^{b}$ ways to choose $B$.
By Lemma 3.3, given $W\cup S\cup B$ and $u$ there are at most $2$ minimal obstructions. Thus, given $u,w,\ell_{2},s,\ell_{3},$ and $b$, the number of minimal obstructions is at most $2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$. ∎
We will further make use of the following estimate on the number of sets whose size is not too small, yet having a small edge-boundary.
Lemma 3.5.
Let $a\geq\frac{n}{d^{\ln^{2}d}}$. Then, the number of sets $A\subseteq V$ of size $a$ with $e(A,A^{C})<a\ln^{2}d$ is at most $\exp\left\{\frac{2a}{\ln d}\right\}$.
Proof.
Let $\mathcal{F}$ be the family of $A\subseteq V$ satisfying the conditions of the lemma.
For $i\in[t]$ and any $A\subseteq V$, let $E_{i}(A,A^{C})\subseteq E(A,A^{C})$ be the set of edges in $E(A,A^{C})$ corresponding to a change in the $i$-th coordinate, and let $e_{i,A}\coloneqq|E_{i}(A,A^{C})|$. Moreover, given $I\subseteq[t]$, let $e_{I,A}=\sum_{i\in I}e_{i,A}$. We say that $A$ is bad with respect to a set of coordinates $I$, if $e_{I,A}<a\ln^{2}d\cdot\frac{|I|}{t}$. Let $\mathcal{A}_{I}$ be the family of sets $A$ which are bad with respect to some $I\subseteq[t]$. Note that for every fixed $m\in\mathbb{N}$, if $A\in\mathcal{F}$, then there is some $I$ with $|I|=m$ such that $e_{I,A}<a\ln^{2}d\cdot\frac{|I|}{t}$. Thus,
$|\mathcal{F}|\leq\sum_{\begin{subarray}{c}I\subseteq[t]\\ |I|=m\end{subarray}}|\mathcal{A}_{I}|\leq\binom{t}{m}\max_{\begin{subarray}{c}I\subseteq[t]\\ |I|=m\end{subarray}}|\mathcal{A}_{I}|.$
We now set $m=\log_{C}\left(\ln^{5}d\right)$, and turn to estimate $|\mathcal{A}_{I}|$ for any $I\subseteq[t]$ with $|I|=m$.
For such an $I$ and $v\in V$, let $G(I,v)\coloneqq\square_{i\in[t]\setminus I}\{v_{i}\}\square_{i\in I}H_{i}\subseteq G$. Observe that $2^{|I|}\leq|V(G(I,v))|\leq C^{|I|}$, and that for every $v\neq u\in V$, $V(G(I,v))$ and $V(G(I,u))$ are either disjoint or identical. Thus, fixing $I$ with $|I|=m$, there are at most $\frac{n}{2^{m}}$ different subgraphs $G(I,v)$, and their union is $V$. We say that $A$ intersects non-trivially with $G(I,v)$ if $V(G(I,v))\cap A\neq\varnothing$ and $V(G(I,v))\setminus A\neq\varnothing$. By Theorem 1.3 and Lemma 3.1, we have that if $A$ intersects non-trivially with $G(I,v)$, then $G(I,v)$ spans at least $|I|=m$ edges of $E(A,A^{C})$. Thus, if $A\in\mathcal{A}_{I}$, we have that $A$ intersects non-trivially at most $\frac{a\ln^{2}d}{t}$ such subgraphs. Indeed, otherwise, there would be no $I$ with $|I|=m$ and $e_{I,A}<a\ln^{2}d\cdot\frac{m}{t}$. Therefore, a set $A\in\mathcal{A}_{I}$ contains at least
$\frac{a-C^{m}\frac{a\ln^{2}d}{t}}{C^{m}}$
such subgraphs, and at most $C^{m}\frac{a\ln^{2}d}{t}$ other vertices. Recalling $a\geq\frac{n}{d^{\ln^{2}d}}$ we thus obtain that
$|\mathcal{A}_{I}|$ $\leq\binom{\frac{n}{2^{m}}}{\frac{a}{C^{m}}}\binom{n}{C^{m}\frac{a\ln^{2}d}{t}}\leq\left(\frac{enC^{m}}{a}\right)^{\frac{a}{C^{m}}}\left(\frac{ent}{C^{m}a\ln^{2}d}\right)^{C^{m}\frac{a\ln^{2}d}{t}}$
$\leq\left(ed^{\ln^{2}d}C^{m}\right)^{\frac{a}{C^{m}}}\left(\frac{etd^{\ln^{2}d}}{C^{m}\ln^{2}d}\right)^{C^{m}\frac{a\ln^{2}d}{t}}\leq\left(ed^{\ln^{2}d}\ln^{5}d\right)^{\frac{a}{\ln^{5}d}}\left(\frac{ed^{\ln^{2}d+1}}{\ln^{7}d}\right)^{\frac{a\ln^{7}d}{t}}$
$\leq\exp\left\{\frac{a}{\ln d}\right\}.$
Altogether, we obtain that $|\mathcal{F}|\leq\binom{t}{\log_{C}(\ln^{5}d)}\exp\left\{\frac{a}{\ln d}\right\}\leq\exp\left\{\frac{2a}{\ln d}\right\}$, as required. ∎
3.3 Proof of Theorem 3b
Recall from Section 1.2 the definition of a minimal obstruction and of a non-trivial minimal obstruction. Further, recall the sets $U,V_{1},W,S$, and $B$, and $\ell,\ell_{1},\ell_{2},\ell_{3}$. Recall that $\omega$ is a function tending arbitrarily slowly to infinity and that $p$ is such that $(1-p)^{d}=\omega/n$.
Before turning to the proof itself, we require some preparation.
For $k\in\{1,\ldots,n\}$, we define
$f(k)=\min_{A\subseteq V(G),|A|=k}e(A,A^{C}).$ (3)
We note that f(k)=f(nk)f(k)=f(n-k), and we may thus restrict our attention to $k\in\{1,\ldots,n/2\}$. Further, by Theorem 1.3,
$f(k)\geq f^{*}(k)\coloneqq\max\left\{k(d-(C-1)\log_{C}k),k\cdot\frac{e}{C}\cdot\ln\left(\frac{n}{k}\right)\right\}.$
Remark 3.6.
We note that, while in general the maximum of two concave functions is not necessarily concave, $f^{*}(k)$ is indeed concave in the domain $\{1,\ldots,n/2\}$ (see, e.g., [12, Section 3.4]). We further note that $f^{*}$ is increasing in $\{1,\ldots,n/e\}$, and decreasing afterwards. Further, $f^{*}(n/2)\geq n/C$.
Given an obstruction $U$, let $W_{1},\dots,W_{w/2}$ denote the components of $G_{p}[W]$, $S_{1},\dots,S_{\ell_{2}}$ denote the components of $G_{p}[S]$ and $B_{1},\dots,B_{\ell_{3}}$ denote the components of $G_{p}[B]$ and set
$\hat{e}\coloneqq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C}).$
We will utilise the following two edge-counting claims. The first concerns $G_{p}[S]$.
Claim 3.7.
Fix an obstruction. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components of $G_{p}[S]$. Then,
$\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})\geq 3(d-1)(\ell_{2}-1)+f^{*}\left(s-3(\ell_{2}-1)\right).$
Proof.
By definition, $\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})\geq\sum_{i=1}^{\ell_{2}}f^{*}(|S_{i}|)$. Further, we have that $\sum_{i=1}^{\ell_{2}}|S_{i}|=s$, and that for each $i\in[\ell_{2}]$, $3\leq|S_{i}|<n/d^{C^{3}/p}$. Thus, by Remark 3.6 and by standard convexity arguments, we have that $\sum_{i=1}^{\ell_{2}}f^{*}(|S_{i}|)$ is minimised when $|S_{i}|=3$ for all $i\in[1,\ell_{2}-1]$, and $|S_{\ell_{2}}|=s-3(\ell_{2}-1)$. Therefore,
$\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})$ $\geq 3(d-1)(\ell_{2}-1)+f^{*}(s-3(\ell_{2}-1)),$
since $f^{*}(3)\geq 3d-3=3(d-1)$. ∎
We abbreviate
$f^{*}_{S}\coloneqq 3(d-1)(\ell_{2}-1)+f^{*}\left(s-3(\ell_{2}-1)\right).$ (4)
The next claim concerns $G_{p}[B]$.
Claim 3.8.
Fix an obstruction. Let $B_{1},\ldots,B_{\ell_{3}}$ be the components of $G_{p}[B]$. Then,
$\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}.$
Proof.
By definition, $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq\sum_{i=1}^{\ell_{3}}f^{*}(|B_{i}|)$. Further, we have that $\sum_{i=1}^{\ell_{3}}|B_{i}|=b$, and that for each $i\in[\ell_{3}]$, $|B_{i}|\geq n/d^{C^{3}/p}$. Thus, by Remark 3.6 and by standard convexity arguments, we have that
$\sum_{i=1}^{\ell_{3}}f^{*}(|B_{i}|)$ $\geq(\ell_{3}-1)f^{*}\left(\frac{n}{d^{C^{3}/p}}\right)+\min\left\{f^{*}\left(\frac{n}{2}\right),f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}$
$\geq(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\cdot\frac{e}{C}\cdot\ln(d^{C^{3}/p})+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}$
$\geq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\},$
as required. ∎
Here we also abbreviate
$f^{*}_{B}\coloneqq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}.$ (5)
The rest of the subsection is structured as follows. First we show that for $u$ of order $n/poly(d)$ the number of components in $B$ is at most one. Together with the trivial bound $\binom{n}{u}$ for the number of obstructions of size $u$, this suffices to show that whp there are no minimal non-trivial obstructions with $u\leq d$ (see Lemma 3.10). Then we turn to obstructions of more substantial size, particularly $d\leq u\leq\frac{n}{d^{C^{5}/p}}$; to handle such obstructions, we first establish efficient enumeration of minimal obstructions of such sizes (Lemma 3.13 and 3.14), and then, utilising these enumerations, show that whp there are no minimal obstructions with such values of $u$ in Lemma 3.15. Then, we turn to obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$. Here, the focus will turn to the set $B$. We will distinguish the cases $b\leq n/2$ and $b>n/2$. If $e(B,B^{C})$ is large, the calculation is straightforward (see Lemma 3.16 and Lemma 3.19). However, if the edge boundary of $B$ is small, the fingerprint argument from Lemma 3.5 comes into play, and allows us to resolve these cases in Lemma 3.17 and Lemma 3.20.
No minimal obstructions with $u\leq\frac{n}{d^{C^{5}/p}}$ and $\ell_{3}\geq 2$.
We begin by showing that for $u\leq\frac{n}{d^{C^{5}/p}}$ whp there are no minimal obstructions such that there are at least two large components in $B$.
Lemma 3.9.
Whp, there are no minimal obstructions with $\ell_{3}\geq 2$ and $u\leq\frac{n}{d^{C^{5}/p}}$.
Proof.
Fix $u$, and fix a minimal obstruction $U$ with $|U|=u$. Denote by $B_{1},\ldots,B_{\ell_{3}}$ the components of $G_{p}[B]$. By Claim 3.8, we have that $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq f^{*}_{B}$. Since $G$ is $d$-regular, $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C}\setminus U)\geq f^{*}_{B}-du$. Note that we may have counted some of the edges twice in the above, and therefore $\left|\bigcup_{i=1}^{\ell_{3}}E(B_{i},B_{i}^{C}\setminus U)\right|\geq(f^{*}_{B}-du)/2$. Note that none of these edges are present in $G_{p}$. There are at most $\binom{n}{u}$ ways to choose $U$, which uniquely defines the obstruction, and thus by the union bound the probability that for a fixed $u$ there is such an obstruction is at most
$\binom{n}{u}(1-p)^{(f^{*}_{B}-du)/2}$ $\leq n^{u}(1-p)^{n\ln d/(3d^{C^{3}/p})}$
$\leq n^{u}\left(\frac{\omega}{n}\right)^{(n\ln d)/(3d^{C^{3}/p+1})}$
$\leq\exp\left\{\ln n\left(u-\frac{n}{4d^{C^{3}/p+1}}\right)\right\}=o(1/n).$
where in the first inequality we used (5) together with our assumption that $\ell_{3}\geq 2$ and $u\leq\frac{n}{d^{C^{5}/p}}$, in the second inequality we used that $(1-p)^{d}=\omega/n$, the penultimate inequality follows since $\omega$ tends arbitrarily slow to infinity, and the last equality follows since we assume $u\leq\frac{n}{d^{C^{5}/p}}$. Union bound over the less than $n$ choices of $u$ completes the proof. ∎
Note that in the above, our choice for the upper-bound on the size of a component in $S$ (and thus a lower-bound on the size of a component in $B$) came into play.
No minimal non-trivial obstructions with $u\leq d$.
We now turn our attention to ‘small’ obstructions.
Lemma 3.10.
Whp, there are no minimal non-trivial obstructions with $u\leq d$.
Proof.
Fix $u$, and fix a minimal non-trivial obstruction. Note that by Lemma 3.9, we may assume that $\ell_{3}\leq 1$. Let $W_{1},\ldots,W_{w/2}$ be the components in $G_{p}[W]$. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components in $G_{p}[S]$. Denote by $v_{1},\ldots,v_{\ell_{1}}$ the vertices of $V_{1}$.
Note that
$\hat{e}\geq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+f^{*}(b).$
Noting that for each $3\leq k\leq\frac{n}{d^{{C^{3}}/p}}$, we have that $f^{*}(k)>k$, we have that
$\hat{e}\geq\ell_{1}d+w(d-1)+s+f^{*}(b).$
Observe that if either $\ell_{1}+w\geq n/10$ or $s\geq n/10$, we would have that the probability of such an obstruction is at most
$\sum_{u=1}^{d}\binom{n}{u}(1-p)^{(\hat{e}-du)/2}\leq dn^{d}\cdot\left(\frac{\omega}{n}\right)^{n/30}=o(1).$
Thus, we may assume that $\ell_{1}+w\leq n/10$ and $s\leq n/10$ (and in particular, $b\neq 0$). In that case, $f^{*}(b)=f^{*}(s+\ell_{1}+w+u)$ is increasing with $s$, and thus our lower bound estimate for $\hat{e}$ is increasing with $s$. We consider two cases separately. First, suppose $s\neq 0$, and thus our estimate is minimised when $s=3$ (and $\ell_{2}=1$). Then $\ell_{1}\geq u-1$ and since our lower bound estimate is also increasing in $\ell_{1}$, our estimate is minimised when $\ell_{1}=u-1$. We then have that
$f^{*}(s+\ell_{1}+w+u)=f^{*}(2u+2)\geq(2u+2)(d-(C-1)\log_{C}(2u+2))\geq(2u+2)(d-2C\ln d).$
Then, the probability of such an obstruction is at most
$\sum_{u=1}^{d}n^{u}(1-p)^{\left(d(u-1)+(2u+2)(d-2C\ln d)-du\right)/2}\leq\sum_{u=1}^{d}n^{u}\left(\frac{\omega}{n}\right)^{u+1/3}\leq d\cdot n^{-1/4}=o(1),$
where we used that $\omega$ tends to infinity arbitrarily slowly.
Now, we turn to the case where $s=0$. In this case, since we consider non-trivial obstructions, we must have that $\ell_{1}\geq u+1$. Then, $\hat{e}\geq(u+1)d+f^{*}(2u+1)\geq(u+1)d+(2u+1)(d-2C\ln d)$. Then, the probability of such an obstruction is at most
$\sum_{u=1}^{d}n^{u}\left(1-p\right)^{((u+1)d+(2u+1)(d-2C\ln d)-du)/2}\leq\sum_{u=1}^{d}n^{u}\left(1-p\right)^{du+d-C\ln d}\leq d\cdot n^{-1/2}=o(1),$
completing the proof. ∎
Note, that in the last part of the proof it was crucial to exclude trivial obstructions.
No minimal obstructions with $d\leq u\leq\frac{n}{d^{C^{5}/p}}$.
We now turn our attention to obstructions of more substantial size.
The following fairly simple observation will be of use throughout the proof.
Claim 3.11.
Suppose $d\leq u\leq\frac{n}{d^{C^{5}/p}}$. Then, whp, there are no obstructions with either $\ell_{1}+w\geq n/10$ or $s\geq n/10$.
Proof.
Similar to the calculation in Lemma 3.10 note that if either $\ell_{1}+w\geq n/10$ or $s\geq n/10$, we would have that the probability of such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}\binom{n}{u}(1-p)^{(\hat{e}-du)/2}\leq\frac{n}{d^{C^{5}/p}}n^{n/d^{C^{5}/p}}\cdot\left(\frac{\omega}{n}\right)^{n/(30d)}=o(1).$
It will also be of use to rule out the existence of such minimal obstructions with $\ell\geq 1.1u$.
Lemma 3.12.
Whp, there are no minimal obstructions with $d\leq u\leq\frac{n}{d^{C^{5}/p}}$ and $\ell\geq 1.1u$.
Proof.
Fix $u$, and fix a minimal obstruction. Note that by Lemma 3.9, we may assume that $\ell_{3}\leq 1$. Let $W_{1},\ldots,W_{w/2}$ be the components in $G_{p}[W]$. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components in $G_{p}[S]$. Denote by $v_{1},\ldots,v_{\ell_{1}}$ the vertices of $V_{1}$.
As before, we note that
$\hat{e}\geq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+f^{*}(b),$
and since for each $3\leq k\leq\frac{n}{d^{{C^{3}}/p}}$, we have that $f^{*}(k)>k$, we obtain
$\hat{e}\geq d\ell_{1}+w(d-1)+s+f^{*}(s+\ell_{1}+w+u).$
By Claim 3.11, we may assume that $\ell_{1}+w\leq n/10$ and $s\leq n/10$ (and in particular, $b\neq 0)$. In that case, $f^{*}(b)=f^{*}(s+\ell_{1}+w+u)$ is increasing with $s,\ell_{1}$ and $w$, and thus our lower bound estimate for $\hat{e}$ is increasing with $s,\ell_{1}$ and $w$. Thus, our estimate is minimised when $s=0$ (and $\ell_{2}=0$), $w=0$ and $\ell_{1}=1.1u$. We then have that
$f^{*}(s+\ell_{1}+w+u)=f^{*}(2.1u)\geq 2.1u(d-C\log_{C}(2.1u))$
Then, the probability of such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}n^{u}(1-p)^{\left(1.1du+2.1u(d-2C\ln u)-du\right)/2}\leq\sum_{u=d}^{n/d^{C^{5}/p}}n^{u}\left(\frac{\omega}{n}\right)^{1.05u}=o(1),$
where we used that $\omega$ tends to infinity arbitrarily slowly and $u\geq d$. ∎
We can now count the number of (relevant) obstructions much more efficiently.
Lemma 3.13.
Suppose $u\geq d$, $\ell\leq 1.1u$ and $\ell_{3}=1$. Then, there are at most
$\binom{n}{\frac{5}{11}u}(ed)^{\frac{6}{11}u+\ell_{1}+s}$
choices for a minimal obstruction.
Proof.
By Lemma 3.2 each vertex $v\in U$ sends edges to at least three components in $G_{p}[V_{1}\cup S\cup B]$, and by our assumptions that $\ell_{3}=1$, it neighbours at least two components in $G_{p}[V_{1}\cup S]$.
Consider now the following bipartite auxiliary graph $F$ with $V(F)=(A_{1},A_{2})$. Set $A_{1}=U$, and let $A_{2}$ be the set of components in $G_{p}[V_{1}\cup S]$. We draw an edge between two vertices $v\in A_{1}$ and $C\in A_{2}$ in this auxiliary graph $F$ if $v$ sends an edge to the component $C$ in $G_{p}$. Note that, by our assumptions, there are at least $2u$ edges in $F$. On the other hand, $|A_{2}|=\ell-1\geq u$ and thus the average degree in $F$ is at least $\frac{2u}{\ell-1}\geq\frac{20}{11}$, using that $d\leq\ell\leq 1.1u$. In a component of size at most three the average degree is at most $\frac{4}{3}$, thus at most $\frac{3}{11}$ of the vertices in $A_{1}$ lie in a component of size at most three (otherwise, the average degree would be less than $\frac{4}{3}\cdot\frac{3}{11}+2\cdot\frac{8}{11}=\frac{20}{11}$).
Going back to the original graph $G$, this remains true, i.e., at least $\frac{8}{11}u$ vertices of $U$ lie in components of size at least four in $G$. Thus, the number of choices for such an obstruction is at most
$\binom{n}{\frac{5}{11}u}(ed)^{\frac{6}{11}u+\ell_{1}+s}.$
In fact, we will use the above counting in order to provide a typical better estimate on the number of minimal obstructions. Here, we will first assume that $s=0$; we will later show that indeed the probability of an obstruction is maximized when $s=0$ (similarly to Lemma 3.12).
Lemma 3.14.
Suppose $d\leq u\leq n/d^{C^{5}/p}$, $\ell\leq 1.1u$, $s=0$ and $\ell_{3}=1$. Then, whp, for every minimal obstruction, there is a $M\subseteq V$ satisfying the following,
  1. 1.
    $U\subseteq M$; and,
  2. 2.
    $|M|\leq 3u$; and,
  3. 3.
    $G[M]$ contains at most $u/(2^{d/(600C)})$ components.
Proof.
By Lemma 3.13, there are at most $\binom{n}{5u/11}(ed)^{\frac{6}{11}u+\ell_{1}+s}$ choices for the obstruction.
In $G$ each vertex in $V_{1}$ sends $d$ edges to other components. Since $U$ is an obstruction in $G_{p}$ all the remaining edges need to lie in $E(U,V_{1})$. Suppose first that $e(U,V_{1})\leq\frac{du}{20}$. Then, we have that
$\hat{e}-e(U,V_{1})\geq d\ell_{1}+f^{*}(b)-du/20\geq d\ell_{1}-du/20.$
Thus, the probability of such an obstruction is at most
$\binom{n}{5u/11}(ed)^{\frac{6}{11}u+\ell_{1}}(1-p)^{(d\ell_{1}-du/20)/2}\leq\left(\frac{11en}{5u}\right)^{5u/11}(ed)^{6u/11+\ell_{1}}\left(\frac{\omega}{n}\right)^{\ell_{1}/2-u/40}.$
Noting that the above is decreasing as $\ell_{1}$ is increasing, and that under our assumptions $\ell_{1}\geq u$, we obtain that the probability of such an obstruction is at most
$\left(\frac{11en}{5u}\right)^{5u/11}(ed)^{17u/11}\left(\frac{\omega}{n}\right)^{19u/40}\leq n^{-u/50},$
where we used that $\omega$ is a function tending arbitrarily slowly to infinity and that $d=\Theta(\ln n)$. Summing over all possible values of $u$, we obtain that whp there is no such obstruction. In what follows, we may thus assume that $e(U,V_{1})>du/20$.
Let $M=U\cup V_{1}$. We have that $U\subseteq M$ and $|M|=u+\ell_{1}\leq u+1.1u\leq 3u$, by our assumption $\ell_{1}<\ell\leq 1.1u$. 1.3 implies that for any set $M\subseteq V(G)$
$e_{G}(M)\leq|M|d-|M|(d-(C-1)\log_{2}|M|)=|M|(C-1)\log_{2}|M|.$
If $G[M]$ had at least $u/2^{\frac{d}{600C}}$ components, then by standard convexity arguments, we would have that
$e_{G}(M)$ $\leq|M|(C-1)\log_{2}\left(\frac{|M|}{u/2^{\frac{d}{600C}}}\right)$
$\leq 3u(C-1)\log_{2}\left(3\cdot 2^{\frac{d}{600C}}\right)<\frac{du}{20},$
a contradiction. ∎
With these lemmas at hand, we are ready to show that whp there are no minimal obstructions for this range of $u$.
Lemma 3.15.
Whp, there are no minimal obstructions with $d\leq u\leq\frac{n}{d^{C^{5}/p}}$.
Proof.
Assume that there is an obstruction with $\sqrt{d}\leq u\leq\frac{n}{d^{C^{5}/p}}$. By Lemma 3.12 and Lemma 3.9, we may assume that $\ell\leq 1.1u$ and $\ell_{3}\leq 1$.
Fix $u$, and fix a such a minimal non-trivial obstruction. Let $W_{1},\ldots,W_{w/2}$ be the components in $G_{p}[W]$. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components in $G_{p}[S]$. Denote by $v_{1},\ldots,v_{\ell_{1}}$ the vertices of $V_{1}$.
We claim that the function
$\binom{n}{5u/11}(ed)^{\frac{6}{11}u+\ell_{1}+s}(1-p)^{(\hat{e}-du)/2}$ (6)
is decreasing as either $w$ or $s$ are increasing, and since (6) is an upper bound on the probability of an obstruction we may assume that $w=0$ and $s=\ell_{2}=0$ (and union bound over the less than $n^{2}$ choices of $s$ and $w$).
Indeed, recall that
$\hat{e}\geq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+f^{*}(b),$
where $\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})\geq w(d-1)$, $\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})\geq f^{*}_{S}$ and $f^{*}(b)\geq f^{*}(\ell_{1}+w+s+u)$. By 3.11 we may assume that $s<n/10$ and $w<n/10$. Then $\ell_{1}+w+s+u\leq n/4$ (recalling that $\ell_{1}\leq\ell\leq 1.1u$) and $f^{*}(b)=f^{*}(\ell_{1}+w+s+u)$ is increasing with $w$ and $s$.
Thus, in order to show that (6) is decreasing as $s$ is increasing, it suffices to observe that $(ed)^{s}(1-p)^{f^{*}_{S}}$ is decreasing when $s$ is increasing (recalling that $f^{*}_{S}\coloneqq 3(d-1)(\ell_{2}-1)+f^{*}(s-3(\ell_{2}-1))$ from (4)). Indeed, increasing the value of $s$ will not force us to increase the value of $\ell_{1}$ (in fact, it may do the opposite), and as we noted $f^{*}(\ell_{1}+w+s+u)$ is increasing with $s$.
Similarly, $(ed)^{w/2}(1-p)^{w(d-1)}$ is decreasing when $w$ is increasing. Thus, since we observed that $f^{*}(\ell_{1}+w+s+u)$ is increasing with $w$, (6) is decreasing as $w$ is increasing.
Assuming $w=s=0$, by Lemma 3.14, we may assume that there exists $M$ such that $|M|\leq 3u$, $U\subseteq M$ and there are at most $u/2^{\frac{d}{(600C)}}\eqqcolon r$ components in $G[M]$. We have at most $n^{r}$ ways to choose the sizes of these components. Thus, by Lemma 1.4, there are at most $\binom{3u}{u}n^{r}(ed)^{3u}$ ways to choose $U$. Thus, the probability of having such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}\binom{3u}{u}n^{r}(ed)^{3u}(1-p)^{(\hat{e}-du)/2},$
where $\hat{e}\geq d\ell_{1}+f^{*}(b)=d\ell_{1}+f^{*}(2u)\geq du+\frac{2eu}{C}\cdot\ln(n/2u)$. We thus have that
$\sum_{u=d}^{n/d^{C^{5}/p}}\binom{3u}{u}n^{r}(ed)^{3u}(1-p)^{(\hat{e}-du)/2}$
$\leq$ $\sum_{u=d}^{n/d^{C^{5}/p}}(3e)^{u}n^{u/2^{\frac{d}{(600C)}}}(ed)^{3u}\exp\left\{-p\cdot\frac{eu}{C}\cdot\ln(n/2u)\right\}$
$\leq$ $\sum_{u=d}^{n/d^{C^{5}/p}}\exp\left\{u\left(\ln(3e)+\frac{\ln n}{2^{d/(600C)}}+3\ln(ed)-\frac{ep}{C}\ln(n/2u)\right)\right\}.$
Recalling that $\ln n=\Theta(d)$ and using that $u\leq n/d^{C^{5/p}}$ the probability of such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}\exp\left\{-uC^{3}\ln d\right\}=o(1/n^{2}),$
where we used that $u\geq d$. Union bound over the less than $n^{2}$ choices of values for $s,w$ completes the proof. ∎
Obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$.
Having treated all obstructions with an upper bound on $u$ we turn to arbitrarily large obstructions. Here we will focus on the set $B$ and distinguish whether $b\leq\frac{n}{2}$ or $b>\frac{n}{2}$.
Before looking at these cases let us remark that again the probability of having an obstruction with given properties is maximised when $s=0$ and $w=0$. Indeed, by Lemma 3.4 the number of obstructions given $u,w,s,\ell_{2},b$ and $\ell_{3}$ is bounded by $2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$. Consider again
$\hat{e}$ $=\ell_{1}d+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})$
$\geq d\ell_{1}+w(d-1)+f^{*}_{S}+\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C}).$
Then, the probability of having such an obstruction is bounded by
$2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}(1-p)^{(\hat{e}-du)/2}.$ (7)
We claim that (7) is decreasing as both $w$ and $s$ are increasing and thus we assume $w=s=0$ in the remaining part of the section (and union bound over the possible values for $w,s$ in the end).
To see this, observe that for each component of size two in $W$, we have
$n(ed)^{2}(1-p)^{2(d-1)}=\frac{(\omega ed)^{2}}{n}(1-p)^{-2}=o(1).$
Similarly, from 3.7 we know that the edge-boundary of $S$ is minimised when all but one component are of size three. For components with three vertices the calculation above holds verbatim, i.e.,
$n(ed)^{3}(1-p)^{3d-3}=\frac{(\omega ed)^{3}}{n^{2}}(1-p)^{-3}=o(1).$
The remaining component of $S$ is of size at most $\frac{n}{d^{C^{3}/p}}$ (by definition of $S$) and for such a component we obtain
$f^{*}\left(\frac{n}{d^{C^{3}/p}}\right)\geq\frac{eC^{2}}{p}\ln d\frac{n}{d^{C^{3}/p}},$
where we used 1.3. Thus, the remaining component of size $s_{max}$ in $S$ expands by a factor of at least $\frac{eC^{2}}{p}\ln d$ and thus we obtain
$n(ed)^{s_{max}}(1-p)^{(eC^{2}/p)\ln d\cdot s_{max}}=\exp(\ln n+s_{max}\ln ed-eC^{2}s_{max}\ln d)=o(1).$
For the remainder of this section fix $u,w,s,\ell_{2},b$, and $\ell_{3}$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $0\neq b=|B|$. Recall that we will bound the probability for $w=s=0$ and take a union bound over the possible values for $w,s$ in the end. Let us start with $b\leq\frac{n}{2}$ and first treat this case under the assumption that the edge boundary of $B$ in $G$ is not too small.
Lemma 3.16.
Whp, there are no obstructions with $0\neq b\leq n/2$ and $e(B,B^{C})\geq b\ln^{2}d$.
Proof.
Similarly to before, we have that
$\hat{e}\geq\ell_{1}d+e(B,B^{C})\geq\ell_{1}d+b\ln^{2}d.$
Now, by Lemma 3.4, there are at most
$2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$
such minimal obstructions. Then, we obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2n^{2\ell_{3}}(ed)^{b}(1-p)^{(\hat{e}-du)/2}$
$\leq 2n^{2\ell_{3}}(ed)^{b}(1-p)^{(\ell_{1}d+n\ln^{2}d-du)/2},$
Using that $(1-p)^{d}=\omega/n$, we obtain
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2n^{2\ell_{3}}(ed)^{s+b}\left(\frac{\omega}{n}\right)^{1/2(\ell_{1}-u)}\exp\left(-\frac{p}{2}b\ln^{2}d\right)$
$\leq 2n^{3/2\ell_{3}+1/2(u-\ell)}\omega^{1/2(\ell_{1}-u)}\exp\left((s+b)\ln(ed)-\frac{p}{2}b\ln^{2}d\right)$
$\leq 2\exp\left(\frac{1}{2}(u-\ell)\ln n+\frac{3}{2}\ell_{3}\ln n+\frac{1}{2}(\ell_{1}-u)\ln\omega+b\ln ed-\frac{p}{2}b\ln^{2}d\right)$
where we used $\ell_{1}+\ell_{3}=\ell$. Since $U$ is an obstruction we have $\ell>u$. Furthermore, note that $\ell_{3}\ll\frac{b}{d^{2}}$ such that $\ell_{3}\ln n\ll\frac{b}{d}$ and recall that $\omega$ tends to infinity arbitrarily slowly. Thus, we obtain
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq\exp\left(b\ln ed-\frac{p}{2}b\ln^{2}d\right)$
$\leq\exp\left(-\Omega\left(d^{2}\ln^{2}d\right)\right),$
where we used that $b\geq\frac{n}{d^{C^{3}/p}}\gg d^{2}$. By the union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$, we have that the probability of an obstruction violating the statement of the claim is at most
$n^{6}\exp\left\{-\Omega\left(d^{2}\ln^{2}d\right)\right\}=o(1).$
We now turn to obstructions where $B$ has a small edge-boundary.
Lemma 3.17.
Whp, there are no obstructions with $0\neq b\leq\frac{n}{2}$ and $e(B,B^{C})\leq b\ln^{2}d$.
Proof.
Note that every set $B\subseteq V$ with $e(B,B^{C})\leq b\ln^{2}d$ is of size at least $\frac{n}{d^{\ln^{2}d}}$ since by 1.3, for $b<\frac{n}{d^{\ln^{2}d}}$,
$e(B,B^{C})\geq\frac{b\ln^{2}d\log_{C}d}{C}>b\ln^{2}d.$
We again have that
$\hat{e}\geq\ell_{1}d+f^{*}(b).$
Furthermore, since we assume $e(B,B^{C})\leq b\ln^{2}d$ and $b\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2b}{\ln d}\right\}$ choices for $B$. Therefore, by Lemma 3.3, there are at most $2\exp\left\{\frac{2b}{\ln d}\right\}$ such minimal obstructions. We obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left(\frac{2b}{\ln d}\right)(1-p)^{(\hat{e}-du)/2}$
$\leq 2\exp\left\{\frac{2b}{\ln d}\right\}(1-p)^{\frac{1}{2}d(\ell_{1}-u)+\frac{f^{*}(b)}{2}}$
$\leq 2\exp\left\{\frac{2b}{\ln d}-\frac{pf^{*}(b)}{2}\right\}\left(\frac{\omega}{n}\right)^{\frac{1}{2}(\ell_{1}-u)}$
$\leq 2\exp\left\{-\frac{1}{2}(\ell_{1}-u)\ln n+\frac{2b}{\ln d}-\frac{pf^{*}(b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}.$
Recall that $\ell=\ell_{1}+\ell_{3}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{-\frac{1}{2}(\ell-u)\ln n+\frac{1}{2}\ell_{3}\ln n+\frac{2b}{\ln d}-\frac{pf^{*}(b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}$
$\leq 2\exp\left\{\frac{3b}{\ln d}-\frac{pf^{*}(b)}{2}\right\}$
where in the last step we used that $u<\ell$, $\ell_{3}\ll\frac{b}{d^{2}}$ and that $\omega$ tends to infinity arbitrarily slowly. Since we assume that $b\leq\frac{n}{2}$ and $n\leq C^{d}$, we have that $\log_{C}b\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f^{*}(b)\geq\frac{b}{C^{2}}$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{b\left(\frac{3}{\ln d}-\frac{p}{2C^{2}}\right)\right\}\leq\exp\left\{-\frac{d^{2}}{3C^{2}}\right\}.$
union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
Now we turn to obstructions with $b>\frac{n}{2}$. The argument is similar to the other case with a focus on $B^{C}$ and again we split into cases depending on the size of the edge-boundary of $B^{C}$. Since in this regime Lemma 3.9 does not hold, we require a more general (and very crude) bound on $\ell_{3}$.
Claim 3.18.
Whp, there are no obstructions such that $\ell_{3}\geq\frac{9u}{d^{2}}\text{ and }u\geq d^{2}.$
Proof.
Suppose that $u\geq d^{2}$ and $\ell_{3}\geq\frac{9u}{d^{2}}$. Then
$\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})$ $\geq f^{*}_{B}\geq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}\geq\left(\frac{9u}{d^{2}}-1\right)\frac{n\ln d}{d^{C^{3}/p}}\geq 8du.$
Recalling that choosing $U$ determines the obstruction, the probability of an obstruction with $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq 8du$ is at most
$\binom{n}{u}(1-p)^{1/2(8du-du)}\leq n^{u}\left(\frac{\omega}{n}\right)^{3.5u}\leq n^{-2.4u},$
where we used $(1-p)^{d}=\omega/n$ for $\omega$ tending to infinity arbitrarily slowly. A union bound over all possible values of $u$, and the at most $n^{2}$ choices for $b$ and $\ell_{3}$ yields that the probability of an event violating the statement of the claim is at most
$3n^{2}\sum_{u=1}^{n}n^{-2.4u}\leq 3n^{2}\cdot 2n^{-2.4}=o(1).$
Note that by Lemma 3.15, we may assume that $u\geq\frac{n}{d^{C^{5}/p}}$, and thus $|B^{C}|\geq\frac{n}{d^{C^{5}/p}}$. We begin by assuming that $B^{C}$ has a large edge-boundary.
Lemma 3.19.
Whp, there are no obstructions with $0\neq b>\frac{n}{2}$ and $e(B^{C},B)\geq|B^{C}|\ln^{2}d$.
Proof.
Again, we consider
$\hat{e}\geq\ell_{1}d+e(B,B^{C})\geq\ell_{1}d+|B^{C}|\ln^{2}d.$
By Claim 3.18, we may assume that $\ell_{3}<\frac{9u}{d^{2}}$, and since $u+1\leq\ell_{1}+\ell_{3}$, we have that
$\hat{e}\geq du-\frac{9u}{d}+|B^{C}|\ln^{2}d\geq du+\frac{u\ln^{2}d}{2}.$
As there are at most $du$ edges touching $U$ in $G$, there are at least $\frac{u\ln^{2}d}{4}$ edges which are not present in $G_{p}$ since we are potentially double-counting any edge between different components outside of $U$. We have at most $\binom{n}{u}$ ways to choose the obstruction. Thus, the probability of having such an obstruction is at most
$\binom{n}{u}(1-p)^{u\ln^{2}d/4}$ $\leq\left(\frac{en}{u}\right)^{u}\exp\left\{-\frac{pu\ln^{2}d}{4}\right\}$
$\leq\exp\left\{u\left(1+\ln n-\ln u-\frac{p\ln^{2}d}{4}\right)\right\}$
$\leq\exp\left\{u\left(1+\frac{C^{5}}{p}\ln d-\frac{p\ln^{2}d}{4}\right)\right\}\leq\exp\left\{-\frac{pu\ln^{2}d}{5}\right\},$
where in the penultimate inequality we used $u\geq\frac{n}{d^{C^{5}/p}}$. Union bound over the at most $n^{6}$ choices of $u,\ell_{1},w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
We now turn the case where $B^{C}$ has a small edge-boundary.
Lemma 3.20.
Whp, there are no obstructions with $0\neq b>\frac{n}{2}$ and $e(B^{C},B)\leq|B^{C}|\ln^{2}d$.
Proof.
Similarly to before, we have that
$\hat{e}\geq\ell_{1}d+e(B,B^{C})\geq\ell_{1}d+f^{*}(n-b).$
Since we assume $e(B,B^{C})\leq|B^{C}|\ln^{2}d$ and $|B^{C}|\geq u\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2|B^{C}|}{\ln d}\right\}=\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ choices for $B^{C}$, and hence for $B$. Therefore, by Lemma 3.3, there are at most $2\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ such minimal obstructions. We thus obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{(\hat{e}-du)/2}$
$\leq 2\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{\frac{1}{2}d(\ell_{1}-u)+\frac{f^{*}(n-b)}{2}}$
$\leq 2\exp\left\{\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}\right\}\left(\frac{\omega}{n}\right)^{\frac{1}{2}(\ell_{1}-u)}$
$\leq 2\exp\left\{-\frac{1}{2}(\ell_{1}-u)\ln n+\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}.$
Recall that $\ell=\ell_{1}+\ell_{3}$ and by Claim 3.18 $\ell_{3}\leq\frac{9u}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{-\frac{1}{2}(\ell-u)\ln n+\frac{1}{2}\ell_{3}\ln n+\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}$
$\leq 2\exp\left\{\frac{9u}{2d^{2}}\ln n+\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}$
$\leq 2\exp\left\{\frac{3(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}\right\},$
where in the penultimate step we used that $u<\ell$ and in the last step we used that $\frac{u}{d^{2}}\ln n\ll\frac{n-b}{\ln d}$ and that $\omega$ tends to infinity arbitrarily slowly. Since $b\geq\frac{n}{2}$, we have that $\log_{C}(n-b)\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f^{*}(n-b)\geq\frac{n-b}{C^{2}}$. Moreover, $n-b\geq u$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{u\left(\frac{3}{\ln d}-\frac{1}{C^{2}}\right)\right\}\leq\exp\left\{-\frac{u}{2C^{2}}\right\}.$
Recalling that $u\geq\frac{n}{d^{C^{5}/p}}$ a union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
We can now conclude our proof.
Proof of Theorem 3b.
By Lemmas 3.10 and 3.15, we have that whp there are no minimal non-trivial obstructions with $u\leq\frac{n}{d^{C^{5}/p}}$. By Lemmas 3.16 and 3.17, we have shown that whp there are no minimal non-trivial obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$ and $b\leq n/2$. By Lemmas 3.19 and 3.20, we have shown that whp there are no minimal non-trivial obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$ and $b\geq n/2$. Thus, we have shown that whp there are no minimal non-trivial obstructions. Thus, by the Tutte-Berge formula, it follows that the giant component in $G_{p}$ has a (nearly-)perfect matching. ∎
4 Discussion
We have extended the classical result concerning hitting times of minimum degree one, connectivity, and the existence of a perfect matching to random subgraphs of regular Cartesian product graphs. In particular, this includes a simplified self-contained version of the connectivity result for bond percolation on the hypercube. Let us mention that, independently, Collares, Doolittle, and Erde use a similar approach – that is, sprinkling with probabilities $p_{1}$ and $p_{2}$ – to show a connectivity result for bond percolation on the permutahedron [8]. There, however, similarly to the approach of [1, 5, 20], one utilises that in $G_{p_{1}}$, large components are relatively well-spread, that is, typically every vertex in $G$ is quite close (in $G$) to a large component of $G_{p_{1}}$. In this paper, we neither require nor utilise such a ‘density’ statement, and instead use the fact that the isolated vertices are ‘sparsely spread’.
We note that Lemma 3.3 does not require anything from the host graph $G$, and Lemma 3.14 only utilises some of the isoperimetric profile of $G$, yet does not utilise the product structure of $G$. Hence, these two results could be of independent interest, in particular in questions concerning the typical existence of a perfect matching under percolation in a more general setting. Still, in this proof, the bound on the number of sets with ‘bad expansion’ from Lemma 3.5 exploits the coordinate structure present in product graphs.
Many other random graph models are known to have typically the same hitting times for minimum degree one, connectivity, and the existence of a perfect matching (see, for example, [15] and the references therein). It is thus natural to ask what are the minimal requirements on $G$ for this phenomenon to hold. As a step towards this, we propose the following question, considering regular graphs with high-degree.
Question 4.1.
Let $G$ be a $d$-regular graph on $n$ vertices, with $d=\omega(1)$ and $n$ divisible by two. What minimal requirements are needed on $G$, such that in the random graph process on $G$, the hitting times for minimum degree one, connectivity, and the existence of a perfect matching are the same?
Utilising the Tutte-Berge formula we aimed for a (nearly-)perfect matching, that is, a matching missing none or only one vertex (in the case when $n$ is odd) of the graph. A natural extension is to look for the threshold to have a matching that covers all but a small fraction of the vertices.
Question 4.2.
Let $G=\square_{i=1}^{t}H_{i}$ for $H_{i}$ connected, regular and of bounded size. What is the threshold $p^{*}$ such that for all $p\geq p^{*}$, whp, the giant component of $G_{p}$ contains a (nearly-)perfect matching?
Acknowledgements
The authors would like to thank Joshua Erde, Mihyun Kang and Michael Krivelevich for their guidance, advice, and fruitful discussions. Special thanks to an anonymous referee for spotting an error in a previous version. The second author was supported in part by the Austrian Science Fund (FWF) [10.55776/W1230].